### v3

# ESG Rating Disagreement and Corporate Total Factor Productivity:Inference and Prediction

Abstract
This paper examines how ESG rating disagreement ($Dis$) affects corporate total factor productivity (TFP) in China based on data of A-share listed companies from 2015 to 2022. We find that $Dis$ reduces TFP, especially in state-owned, non-capital-intensive, and low-pollution firms. Mechanism analysis shows that green innovation strengthens the dampening effect of $Dis$ on TFP, and that $Dis$ lowers corporate TFP by increasing financing constraints. Furthermore, XGBoost regression demonstrates that $Dis$ plays a significant role in predicting TFP, with SHAP showing that the dampening effect of ESG rating disagreement on TFP is still pronounced in firms with large $Dis$ values.
JEL Codes: G32, G24, C54.
Keywords: ESG rating disagreement; Total factor productivity; Green innovation; Financing constraints; Machine learning
1 Introduction
Since 1987, the United Nations has reinforced its commitment to sustainable development, and introduced the concept of ESG (Environmental, Social, and Governance) by then-UN Secretary-General Kofi Annan in 2004. Since then, ESG principles have been increasingly integrated into global investment strategies and corporate assessments, with their specific meanings evolving over time. In recent years, China is prominently advocating the philosophy that "green mountains and clear waters are as valuable as mountains of gold and silver", and promoting sustainable development both domestically and internationally. On December 27, 2023, the Guidelines for the Comprehensive Advancement of Beautiful China Construction was issued by the Central Committee of the Communist Party of China and the State Council clearly clarified the exploration and development of ESG assessment frameworks, which shows a strong will of establishing ESG Evaluation System with Chinese Characteristics.
Despite the implementation of various ESG policies in recent year, there are still significant disagreement in ESG rating system. From the moment the ESG concept was introduced in China, multiple rating agents have independently published their ESG rating reports. However, these reports often display significant discrepancies in ratings, which undoubtedly has a substantial impact on the companies being evaluated. Examining the economic consequences of these ESG rating disagreements is of considerable value for both enterprises and regulatory bodies. The main recent studies are: Wang et al. (2024); Zhao et al. (2024) argue that disagreements in ESG ratings are significantly negatively correlated with stock returns; ESG rating disagreement increase analyst forecast errors(Liu et al., 2024); ESG rating disagreement drives firms towards green innovation (Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); Moreover, increased disagreement in ESG ratings leads to higher levels of real earnings management both in the current period and over the next one to two years (Li and Cheng, 2024); The firms with agreed high ESG rating can benefit through cost reduction and larger amount debt-financing(Guo et al., 2024); Besides,the disagreement in ESG ratings may poses a potential risk of stock price collapse(Luo et al., 2023); What’s more, ESG rating disagreement significantly increases audit fees due to information asymmetry(Ling et al., 2024).
However, the focus of most recent studies has been less on exploring the impact of ESG rating disagreements on the overall development of a company. Therefore, we select TFP as the subject of our analysis in this study, as it serves as an indicator of a company’s high-quality development.
2 Theoretical Analysis and Research Hypotheses
According to information asymmetry theory, ESG rating disagreements can confuse stakeholders and contribute to the degree of information asymmetry between companies and investors. ESG ratings, as non-financial information, provide stakeholders with a more comprehensive insight into companies, thereby alleviating information asymmetry(Xiaohong and Zhenghan, 2023). Conversely, significant ESG rating disagreements can intensify information asymmetry, creating a noise effect that leads to investor misjudgments(Berg et al., 2022). Thereby, increasing financing constraints and ultimately resulting in a decline in total factor productivity. Simultaneously, based on stakeholder theory, companies may take certain remedial measures in response to ESG rating disagreements to balance the interests of various stakeholders. For example, companies might intensify green innovation under the stimulus of ESG rating disagreements(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); However, as this initiative cannot cater to all stakeholders under noise effect caused by ESG rating disagreement, it may lead some groups to perceive "greenwashing" behaviors, which ultimately undermines the company’s financing efforts and results in a decline in total factor productivity(Peng and Xie, 2024).
In summary, we propose the research hypothesis H1: ESG rating disagreement reduces corporate TFP.
3 Data and Empirical Design
3.1 Sample Selection
We utilizes financial data of A-share listed companies in China from 2015 to 2022111Before 2015, the only more well-known local ESG rating agencies in China were Huazheng, alongside ESG ratings from four agencies: Huazheng, Wind, SynTao Green Finance, and MioTech. The data processing procedures follow these steps: (1) excluding companies in the financial and real estate industries; (2) excluding ST, *ST, and delisted companies; (3) excluding companies with only one ESG rating in a given year; (4) excluding companies with missing key indicators; and (5) applying Winsorization at the 1% level on both tails. All data is sourced from the CSMAR database, the CNRDS database, and the four rating agencies, resulting in 13417 samples in baseline regression.
3.2 Model Specification
To test the hypotheses presented earlier,we conduct the following baseline regression model [^2] Considering the time required for information to be transmitted and for the market to react, TFP is lagged by one period:
$TFP_{i,t+1}=\alpha_{0}+\alpha_{1}Dis_{i,t}+\alpha_{2}Control_{i,t}+Year_{t}+Id_{i}+\epsilon_{i,t}$ (1)
where $TFP_{i,t+1}$ is the total factor productivity of company $i$ in year $t+1$, $Dis_{i,t}$ is the ESG rating discrepancy, $Control_{i,t}$ represents the control variables, $Year_{t}$ and $Id_{i}$ denote year and company fixed effects, respectively, and $\epsilon_{i,t}$ is the error term.
3.3 Variable Descriptions
(1) Dependent Variable
The dependent variable in this study is Corporate Total Factor Productivity. Considering the robustness of the empirical process, we use three methods to measure TFP, they are LP introduced byLevinsohn and Petrin (2003) and adopted by Lu and Lian (2012), OP by Olley and Pakes (1992) and GMM by Blundell and Bond (1998). In the prediction task, we use TFP measured by LP method as the dependent variable(Xue et al., 2024).
(2) Independent Variable
The independent variable in this study is ESG Rating Disagreement. Domestic ESG rating agencies have advantages in terms of time and spatial proximity that foreign agencies cannot match, making them more likely to influence the behavior of domestic firms. Therefore, this study uses ESG rating data from four Chinese agencies: Huazheng, Wind, SynTao Green Finance, and MioTech. Following Avramov et al. (2022), the ratings were assigned positive values according to their rating levels, with a difference of 1 between levels. We define the rankings as the values normalized to the range [0, 1] for the given rating agency in a given year, and the pairwise standard deviation of these rankings is used as the measure of $Dis$.
$Dis_{i,t}=\sqrt{\frac{1}{n-1}\sum_{k\neq j}^{n}(rank_{i,t,k}-rank_{i,t,j})^{2}}$ (2)
where $rank_{i,j,k}$ represents the normalized ranking of corporate $i$ in year $t$ given by agency $k$, $n$ represents the number of pairs, $k$ and $j$ represent different ESG rating agencies.
(3) Control Variables
Based on the characteristics of our research and with reference to the works of Yang et al. (2024) and Xue et al. (2024), the control variables include firm value, top 10 shareholders’ ownership percentage, years listed, leverage ratio, firm size, revenue per capita, and ownership ratio. The main variables are defined in Table 1.
Table 1: Definition of Main Variables
Symbol Source Description
$TFP$ CSAMR Estimated using the LP, OP and GMM method
$Dis$ ESG Rating Agencies Calculated based on the method described above
$TobinQ$ CSAMR Tobin’s Q ratio
$Top10$ CSAMR Shareholding of top 10 shareholders / total shares
$ListAge$ CSAMR Natural logarithm of years since the firm’s IPO
$Lev$ CSAMR Total liabilities at year-end / total assets at year-end
$Size$ CSAMR Natural logarithm of total assets at year-end
$Avg$ CSAMR Natural logarithm of operating income / number of employees
$Der$ CSAMR Total liabilities at year-end / total equity at year-end
$EI$ CNRDS Natural logarithm of the number of green patents obtained in a given year
$FC$ CSAMR WW and KZ indices, the larger they are the more constrained the financing is
4 Empirical Analysis
Table 2: Descriptive Statistics of Main Variables
Variables Count Mean SD Min 25% 50% 75% Max
$TFP_{LP}$ 13417 8.51 1.04 6.38 7.79 8.40 9.14 11.28
$TFP_{OP}$ 13417 6.83 0.85 5.14 6.24 6.72 7.33 9.17
$TFP_{GMM}$ 13417 5.72 0.80 4.09 5.18 5.63 6.18 8.12
$Dis$ 13417 0.26 0.19 0.01 0.11 0.24 0.34 0.83
$TobinQ$ 13417 2.03 1.32 0.83 1.22 1.61 2.31 8.34
$Top10$ 13417 56.77 15.15 23.59 45.91 56.95 67.87 90.89
$ListAge$ 13417 2.34 0.68 1.10 1.79 2.40 3.00 3.37
$Lev$ 13417 0.43 0.19 0.06 0.28 0.43 0.57 0.88
$Size$ 13417 22.52 1.34 20.09 21.54 22.32 23.31 26.41
$Avg$ 13417 4.79 0.80 3.07 4.25 4.69 5.25 7.17
$Der$ 13417 1.05 1.07 0.07 0.39 0.74 1.31 6.75
$EI$ 13417 1.30 1.49 0.00 0.00 1.10 2.30 7.74
$WW$ 10059 -1.03 0.08 -1.24 -1.08 -1.03 -0.98 -0.86
$KZ$ 13417 1.17 1.94 -4.35 0.02 1.36 2.49 6.22
  • Note: The data in the table are descriptive statistics for deleting missing values after lagging one period of TFP.
4.1 Baseline Regression
Table 3 reports the results of the baseline regression, with clustering at the individual and time levels. The results indicate that ESG rating disagreement reduces corporate total factor productivity, providing preliminary support for the validity of Hypothesis H1.
Table 3: Baseline Regression Results
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0675*** -0.0536** -0.0536** -0.0447*** -0.0357** -0.0373**
(-2.7452) (-2.1838) (-2.2098) (-2.9449) (-2.2503) (-2.2395)
$TobinQ$0.0032 -0.0017 -0.0007
(-0.5471) (-0.3953) (-0.1795)
$Top10$0.0010 -0.0009 -0.0005
(-0.9976) (-0.9496) (-0.4838)
$ListAge$ 0.0989** 0.0855** 0.0160
(2.4952) (2.3667) (0.4241)
$Lev$0.1697 -0.1557 -0.1481
(-1.5121) (-1.5361) (-1.4348)
$Size$ 0.3737*** 0.2581*** 0.2245***
(13.129) (14.214) (10.939)
$Avg$ 0.2292*** 0.2695*** 0.2671***
(2.7852) (2.8274) (2.7426)
$Der$$0*$0190* -0.0211**
(-2.0405) (-1.8342) (-1.9816)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Doubel Clustering YES YES YES YES YES YES
$N$ 13417 13417 13417 13417 13417 13417
$R^{2}$ 0.0015 0.0011 0.0011 0.1785 0.1670 0.1409
  • Note: t-statistics are in parentheses. * p<0.1, ** p<0.05, *** p<0.01. The same applies below.
4.2 Robustness Tests
To ensure the robustness of the baseline regression results, the following robustness tests were conducted.
First, to verify whether Hypothesis H1 holds when only leading $Dis$ by one period, we lead the it by one period(see table8).
Second, since the total factor productivity of firms in this study is influenced by regional productivity levels, to mitigate endogeneity issues arising from omitted variable bias, we include the provincial-level New quality productivity measured by Lu et al. (2024) as a control variable (see table9).
Third, following Yang et al. (2022), we calculate the industry average of ESG rating disagreements by year and industry, and use the cube of the difference between the industry average and the ESG rating disagreements as an instrumental variable. A two-stage least squares (2SLS) estimation is then conducted to further mitigate potential endogeneity issues (see table10).
Fourth, we denote disagreements under 20% as non-existent and the remainder as present. Propensity scores are computed via logistic regression, followed by 1:1 nearest-neighbor matching and subsequent regression (see table11).
4.3 Mechanism Analysis
4.3.1 The Role of Green Innovation
Myriad existing researches demonstrate that ESG rating disagreement can drive firms to enhance green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b). However, does this green innovation truly contribute to substantial corporate development? This study introduces green innovation as a mechanism variable, measured by the natural logarithm of the number of green patents held by a company. To examine the role of green innovation in the relationship between ESG rating disagreement and TFP, an interaction term between green innovation and ESG rating disagreement is constructed, with other settings remaining consistent with Equation 1.
Table 4 and figure1 reports the results. The results show that when there are no ESG rating discrepancies, green innovation promotes the improvement of TFP. Conversely, when there are ESG rating discrepancies, it strengthens the dampening effect of ESG rating disagreement on TFP. Dose this contradicts exiting researches that show firms tend to increase their green innovation efforts when ESG rating disagreement is present(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b). This observation is very likely caused by a remedial action taken by firms in response to ESG rating disagreement, as Hou and Xie (2024) mentioned, such ESG rating disagreement is negatively correlated with the quality of corporate green innovation. According to stakeholder theory, firm’s remedial green innovations may be seen as "greenwashing" when there is disagreement over ESG ratings, which will lead to higher costs of debt financing for corporations, ultimately bringing about a decline in TFP(Peng and Xie, 2024).
Table 4: Green Innovation Mechanism
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0264 -0.0205 -0.0199 -0.0167 -0.0130 -0.0136
(-1.2045) (-1.0114) (-0.9590) (-0.8367) (-0.7014) (-0.7069)
$EI$ 0.0412*** 0.0283*** 0.0253*** 0.0223*** 0.0142*** 0.0126***
(8.4064) (6.2383) (5.4522) (4.9596) (3.3848) (2.8967)
$Dis*EI$0.0366*** -0.0292*** -0.0296*** -0.0249** -0.0199** -0.0207**
(-3.1706) (-2.7348) (-2.7184) (-2.3693) (-2.0401) (-2.0432)
$TobinQ$0.0030 -0.0016 -0.0007
(-0.7895) (-0.4611) (-0.1778)
$Top10$$0$$0$0005
(-1.7413) (-1.7832) (-0.8682)
$ListAge$ 0.1012*** 0.0869*** 0.0172
(3.4856) (3.2139) (0.6122)
$Lev$0.1692*** -0.1550*** -0.1472***
(-3.2834) (-3.2318) (-2.9594)
$Size$ 0.3664*** 0.2537*** 0.2208***
(28.243) (21.017) (17.632)
$Avg$ 0.2297*** 0.2698*** 0.2673***
(22.666) (28.608) (27.331)
$Der$$0**$0190*** -0.0211***
(-3.1573) (-3.0646) (-3.2805)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
$N$ 13417 13417 13417 13417 13417 13417
$R^{2}$ 0.0092 0.0052 0.0042 0.1806 0.1680 0.1417
Figure 1: Marginal Effect
4.3.2 The Role of Financing Constraints
Undoubtedly, ESG investment plays a crucial role in promoting ESG development, making it essential to examine the impact of ESG rating disagreement on corporate financing. According to the information asymmetry theory discussed in the previous section, ESG rating disagreement introduces noise, potentially contributes financing constraints for firms. This study focuses exclusively on data from domestic ESG rating agencies in China, in contrast to Li et al. (2024), which also includes data from foreign rating agencies to explore this mechanism. Hence, we employ the WW and KZ indices form CSAMR to enhance the robustness of the empirical evidence. Furthermore, we follow Jiang (2022)’s research on the transmission mechanism, to avoid financing constraints being both exogenous and endogenous. We only use independent variable to regress with Finacing constraints, since numerous researches have demonstrated that financing constraints are detrimental to TFP improvement(Hopenhayn, 2014; Zhang et al., 2021; Piao et al., 2023).
Table 5 reports the regression results. The $Disb$ is the variable by one period. So we concludes that increased ESG rating disagreement reduces corporate financing constraints, thereby lowering TFP.
Table 5: Financing Constraints Mechanism
Variables WW
$WW$ $WW$ $KZ$ $KZ$ $WW$ $WW$ $KZ$ $KZ$
$Dis$ 0.0122*** 0.2166*** 0.0089*** 0.1815***
(4.7408) (3.0974) (3.6730) (2.6598)
$Disb$ 0.0122*** 0.2142*** 0.0074*** 0.1392**
(4.7691) (3.0604) (3.3430) (2.2547)
$TobinQ$0.0022*** -0.0006 0.2724*** 0.0237
(-3.9269) -0.9530 (19.528) (1.4433)
$Top10$0.0007*** -0.0006*** 0.0011 -0.0061**
(-8.8614) -6.9840 (0.4783) (-2.3496)
$ListAge$ 0.0204*** 0.0143*** 0.5530*** 1.2048***
(6.1906) (2.9544) (6.5333) (9.5882)
$Lev$ 0.0370*** 0.0359*** 6.7446*** 3.1453***
(5.0820) (4.5416) (32.842) (14.109)
$Size$0.0614*** -0.0455*** -0.4512*** -0.2274***
(-36.832) (-22.973) (-9.5683) (-4.0781)
$Avg$$0**$0113*** -0.5288*** -0.0107
(-18.835) (-7.3282) (-13.531) (-0.2432)
$Der$ 0.0030*** 0.0011 0.0191 -0.0027
(3.4646) (1.2994) (0.6820) (-0.0931)
Controls NO NO NO YES YES YES YES YES
Year YES YES YES YES YES YES YES YES
Id YES YES YES YES YES YES YES YES
$N$ 10059 10059 13417 13417 10059 10059 13417 13417
$R^{2}$ 0.0034 0.0034 0.0010 0.0010 0.2647 0.1221 0.2250 0.0517
  • Note: Similar to the robustness tests, $Dis$ corresponds to regressions on ESG rating disagreement in year t and financing constraints in year t+1, while $Disb$ corresponds to regressions on ESG rating disagreement in year t-1 and financing constraints in year t.
4.4 Heterogeneity Analysis
To further explore whether the dampening effect of ESG rating disagreement on TFP exhibits heterogeneity, this study conducts a subgroup regression analysis. Given that the sample size decreases after subgrouping, making the results more susceptible to the influence of outliers, the data was Winsorized at the 2% level on both ends before performing the subgroup regression and conducting a Chow Test on $Dis$.
Table 6 reports the results of the heterogeneity analysis. The results indicate that the effect is more significant in state-owned firms, non-capital-intensive firms, and low-pollution firms. The above heterogeneity may exist because, first, there is a greater willingness to engage in ESG construction among non-state-owned firms due to the fact that they tend to focus more on ESG performance to attract more investment; second, usually capital-intensive firms are less affected by market forces due to their own low variable costs and high fixed costs; third, in the case of highly polluting firms that are susceptible to a variety of environmental policies and regulations in their financing, such firms are less likely to invest in ESG performance, and therefore the impact of divergent ESG ratings is less significant.
Table 6: Heterogeneity Analysis
Variables TFP
State-Owned Non-State-Owned Capital-Intensive Non-Capital-Intensive High-Pollution Low-Pollution
$TFP_{LP}$
$Dis$0.0622** -$0$0322 -0.0462*** -0.0239 -0.0481***
(-2.0516) (-1.8294) (-1.1009) (-2.8884) (-1.0401) (-2.7959)
Chow Test 500.19*** 289.12*** 544.57***
$R^{2}$ 0.1573 0.1886 0.1741 0.1817 0.2010 0.1754
$TFP_{OP}$
$Dis$0.0553* -0.0268 -0.0097 -0.0458*** -0.0070 -0.0493**
(-1.8608) (-1.5208) (-0.3229) (-2.6338) (-0.3149) (-2.4735)
Chow Test 529.58*** 269.31*** 534.66***
$R^{2}$ 0.1759 0.1593 0.1672 0.1663 0.1930 0.1595
$TFP_{GMM}$
$Dis$0.0618** -$0$0278 -0.0471*** -0.0134 -0.0519**
(-2.1381) (-1.6745) (-0.9906) (-2.6889) (-0.6185) (-2.5484)
Chow Test 498.97*** 248.15*** 475.42***
$R^{2}$ 0.1607 0.1254 0.1504 0.1377 0.1691 0.1321
Controls YES YES YES YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Double Clustering YES YES YES YES YES YES
$N$ 4416 9001 2342 11075 3900 9517
5 Further Discussion
In the above analysis we can get that ESG rating disagreement brings about a decrease in TFP, which is obtained based on linear regression. In order to further explore the relationship between ESG rating divergence and TFP under the nonlinear assumption, we consider a regression analysis using a machine learning approach that can be interpreted with SHapley Additive exPlanations(SHAP) introduced by Scott et al. (2017), which can observe whether ESG rating disagreement is crucial for explaining TFP under the assumption of nonlinear system and explore the nonlinear influence relationship between the two variables.
Following Xue et al. (2024), we use the XGBoost machine learning model introduced by Chen and Guestrin (2016) to predict $TFP_{LP}$. XGBoost is an ensemble learner based on gradient boosted trees (GBDT) that has demonstrated excellent performance in many applications, even outperforming deep neural networks in some tasks. In this paper, the Optuna library developed by Akiba et al. (2019), combed with early stopping and cross-validation strategies is employed for parameter selection. After that, we adopt the SHAP (SHapley Additive exPlanations) method proposed by Scott et al. (2017) to explain XGBoost. The variables include the ESG rating disagreement ($Dis$ $Disb$), all control variables, and institutional ownership percentage ($Inst$).
We use 80% of the dataset for training and 20% for testing and conducted three rounds of training: the first round included ESG rating disagreement, the second round added ESG rating disagreement from the previous period, and the third round did not include ESG rating disagreement. Table 7 reports the evaluation metrics for the three rounds of training. It can be observed that using $R^{2}$ as the evaluation metric, the model’s predictive performance improved by 4.3% and 6.23% respectively compared to the model without ESG rating disagreement, indicating that ESG rating disagreement aids in the short-term prediction of $TFP_{LP}$.
Table 7: Performance Metrics for Training and Test Sets
Metric Training Set 1 Training Set 2 Training Set 3 Test Set 1 Test Set 2 Test Set 3
MSE 0.2609 0.2274 0.2818 0.3492 0.3300 0.3835
RMSE 0.5108 0.4769 0.5309 0.5909 0.5744 00.6192
MAE 0.3995 0.3729 0.4145 0.4589 0.4523 0.4818
MAPE 4.7786% 4.4569% 4.9457% 5.4649% 5.3644% 5.7293%
R2 0.7685 0.7969 0.7500 0.7045 0.7176 0.6755
To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in the predictive task, SHAP was used to interpret the XGBoost model. SHAP values, rooted in game theory, quantify each feature’s contribution to the model output by treating features as "players" in a game. The Shapley value $\phi_{i}$ for a feature $i$ is calculated as:
$\phi_{i}=\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!\cdot(|N|-|S|-1)!}{|N|!}\left[f(S\cup\{i\})-f(S)\right]$ (3)
where $S$ is a subset excluding $i$, $f(S)$ is the model output for $S$, and $f(S\cup\{i\})$ is the output after adding $i$ to $S$.
In this study, based on the previous training results, we chose the model with one-period lagged ESG rating disagreement for our SHAP analysis, and the SHAP values were calculated for the test set. Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting total factor productivity. Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s total factor productivity, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task. The reason for this phenomenon is that the investment market is more tolerant of small ESG rating divergence, and listed companies will not face significant financing pressure due to small divergence.
Figure 2: Feature Importance
Figure 3: SHAP Beeswarm Plot
6 Conclusion
ESG rating disagreement poses challenges to the enhancement of corporate TFP. This study utilizes financial data from China’s A-share listed companies and ESG rating information from domestic ESG rating agencies between 2015 and 2022 to explore the relationship between ESG rating disagreements and firms’ TFP. The findings robustly indicate that ESG rating disagreements have a negative impact on firms’ TFP. The mechanism analysis reveals that green innovation instead strengthens the dampening effect of ESG rating disagreement on TFP; moreover, ESG rating disagreement strengthens firms’ financing constraints, thereby hindering the improvement of TFP. Heterogeneity analysis further indicates that the negative impact of ESG rating disagreements on TFP is more pronounced in state-owned enterprises, non-capital-intensive firms, and companies with lower levels of pollution. Additionally, an important finding from the analysis using machine learning methods suggests that ESG rating disagreements serve as a significant indicator for short-term predicting firms’ TFP, providing new insights for companies to engage in ESG risk management to enhance TFP.
Appendix A Appendix
A.1 Robustness Tests
Table 8: Robustness Tests (I)
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0662*** -0.0524** -0.0524** -0.0246** -0.0159** -0.0180**
(-2.7472) (-2.2660) (-2.2781) (-2.2085) (-2.4677) (-2.0769)
$TobinQ$ 0.0218*** 0.0141*** 0.0150***
(6.8126) (7.0455) (6.1013)
$Top10$ 0.0005 0.0002 0.0004
(0.8363) (0.5950) (0.8865)
$ListAge$$0*$0291** -0.0678***
(-2.3362) (-2.2563) (-4.0705)
$Lev$ 0.1415 0.0739 0.0650
(1.8712) (1.5504) (1.2862)
$Size$ 0.4212*** 0.2448*** 0.1800***
(21.576) (18.539) (11.207)
$Avg$ 0.7135*** 0.8172*** 0.8193***
(38.466) (64.699) (55.668)
$Der$0.0257** -0.0113 -0.0135
(-2.0169) (-1.3877) (-1.5955)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Doubel Clustering YES YES YES YES YES YES
$N$ 13417 13417 13417 13417 13417 13417
$R^{2}$ 0.0015 0.0011 0.0011 0.7027 0.8496 0.7734
Table 9: Robustness Tests (II)
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0647** -0.0519** -$0*$0419*** -0.0342** -0.0346**
(-2.5465) (-1.9984) (-2.0263) (-2.6444) (-1.9838) (-1.9658)
$TobinQ$0.0027 -0.0013 -0.0003
(-0.4030) (-0.2619) (-0.0725)
$Top10$0.0009 -0.0009 -0.0005
(-0.9777) (-0.9697) (-0.5039)
$ListAge$ 0.0906** 0.0796** 0.0058
(2.4895) (2.3720) (0.1639)
$Lev$0.1555 -0.1408 -0.1334
(-1.3079) (-1.2639) (-1.1864)
$Size$ 0.3762*** 0.2596*** 0.2262***
(14.285) (13.442) (10.036)
$Avg$ 0.2277*** 0.2680** 0.2660**
(2.6951) (2.7597) (2.6807)
$Der$$0$0194 -0.0217
(-1.9537) (-1.7844) (-1.9273)
$N.TFP_{province}$ 0.0993 0.0847 0.0403
(0.6302) (0.5857) (0.2531)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Doubel Clustering YES YES YES YES YES YES
$N$ 12893 12893 12893 12893 12893 12893
$R^{2}$ 0.0014 0.0011 0.0010 0.1793 0.1670 0.1406
Table 10: Robustness Tests (III)
Variables Dependent Variable
$Dis$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$IV(\hat{Dis})$0.3336*** -0.0479*** -0.0408** -0.0424**
(-602.14) (-3.0748) (-2.3966) (-2.3533)
$TobinQ$0.0007 -0.0027 -0.0013 -0.0004
(-0.5971) (-0.4037) (-0.2634) (-0.0756)
$Top10$ 3.476e-05 -0.0010 -0.0009 -0.0005
(0.3084) (-0.9808) (-0.9708) (-0.5052)
$ListAge$0.0056 0.0910** 0.0799** 0.0061
(-0.9005) (2.4981) (2.3788) (0.1734)
$Lev$0.0029 -0.1554 -0.1407 -0.1333
(-0.3495) (-1.3058) (-1.2618) (-1.1844)
$Size$0.0045** 0.3760*** 0.2594*** 0.2260***
(-2.6017) (14.301) (13.453) (10.029)
$Avg$0.0017 0.2278*** 0.2680** 0.2660**
(-0.7314) (2.6949) (2.7600) (2.6811)
$Der$ 0.0037** -$0$0193 -0.0216
(2.8094) (-1.9412) (-1.7701) (-1.9119)
$N.TFP_{province}$ 0.0100 0.0999 0.0851 0.0408
(0.5583) (0.6350) (0.5902) (0.2565)
Controls YES YES YES YES
Year YES YES YES YES
ID YES YES YES YES
Double Clustering YES YES YES YES
$F$ 2.983e+04 227.22 208.52 170.30
$N$ 16907 12892 12892 12892
R2 0.9542 0.1794 0.1671 0.1408
  • Note: The second stage of 2SLS regression follows the baseline regression with a one-period lag.
Table 11: Robustness Tests (IV)
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.1054*** -0.0959*** -0.0908*** -$0*$0676** -0.0645**
(-3.0359) (-2.9849) (-2.7841) (-2.3040) (-2.2900) (-2.1303)
$TobinQ$ 0.0007 0.0002 -0.0026
(0.0934) (0.0351) (-0.3889)
$Top10$ 0.0005 0.0004 0.0007
(0.5206) (0.3886) (0.6853)
$ListAge$ 0.1119 0.1001 0.0149
(1.6127) (1.5470) (0.2245)
$Lev$0.0295 -0.0270 -0.0174
(-0.3276) (-0.3216) (-0.2021)
$Size$ 0.3899*** 0.2685*** 0.2291***
(16.968) (12.538) (10.440)
$Avg$ 0.2343*** 0.2761*** 0.2832***
(13.111) (16.582) (16.598)
$Der$0.0408*** -0.0368*** -0.0385***
(-3.4013) (-3.2928) (-3.3577)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
ID YES YES YES YES YES YES
Double Clustering NO NO NO NO NO NO
$N$ 5368 5368 5368 5368 5368 5368
R2 0.0028 0.0027 0.0024 0.1739 0.1628 0.1455
A.2 Optuna Parameter Range & Correlation Coefficient
Table 12: Correlation Matrix
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $Dis$ $TobinQ$ $Top10$ $ListAge$ $Lev$ $Size$ $Avg$ $Der$ $N.TFP_{province}$ $IV$
$TFP_{LP}$ 1.000 0.951 0.918 -0.066 -0.200 0.127 0.329 0.444 0.801 0.648 0.344 -0.037 0.082
$TFP_{OP}$ 0.951 1.000 0.976 -0.054 -0.183 0.104 0.303 0.405 0.707 0.841 0.325 -0.036 0.067
$TFP_{GMM}$ 0.918 0.976 1.000 -0.041 -0.132 0.080 0.244 0.356 0.583 0.823 0.295 -0.005 0.059
$Dis$0.066 -0.054 -0.041 1.000 0.014 -0.078 0.025 0.039 -0.071 -0.024 0.052 -0.015 -0.974
$TobinQ$0.200 -0.183 -0.132 0.014 1.000 -0.036 -0.094 -0.249 -0.283 -0.116 -0.173 0.012 -0.008
$Top10$ 0.127 0.104 0.080 -0.078 -0.036 1.000 -0.364 -0.066 0.147 0.044 -0.052 0.038 0.087
$ListAge$ 0.329 0.303 0.244 0.025 -0.094 -0.364 1.000 0.292 0.424 0.204 0.240 -0.184 -0.024
$Lev$ 0.444 0.405 0.356 0.039 -0.249 -0.066 0.292 1.000 0.458 0.256 0.843 -0.044 -0.038
$Size$ 0.801 0.707 0.583 -0.071 -0.283 0.147 0.424 0.458 1.000 0.413 0.359 -0.104 0.082
$Avg$ 0.648 0.841 0.823 -0.024 -0.116 0.044 0.204 0.256 0.413 1.000 0.222 -0.041 0.027
$Der$ 0.344 0.325 0.295 0.052 -0.173 -0.052 0.240 0.843 0.359 0.222 1.000 -0.047 -0.046
$N.TFP_{province}$0.037 -0.036 -0.005 -0.015 0.012 0.038 -0.184 -0.044 -0.104 -0.041 -0.047 1.000 0.014
$IV$ 0.082 0.067 0.059 -0.974 -0.008 0.087 -0.024 -0.038 0.082 0.027 -0.046 0.014 1.000
Table 13: Hyperparameters for XGBoost Model
Parameter Range(optuna) / Value
loss_function rmse
learning_rate [0.005, 0.2] (log scale)
max_depth [1, 3]
min_child_weight [1, 5] (log scale)
subsample [0.5, 0.8]
colsample_bytree [0.5, 0.8]
reg_lambda [1, 10] (log scale)
reg_alpha [1, 10] (log scale)
tree_method hist
device cuda
num_boost_round 1000
nfold 20
early_stopping_rounds 5
seed 42

### v4

@@
 This paper examines how ESG rating disagreement ($Dis$) affects corporate total factor productivity (TFP) in China based on data of A-share listed companies from 2015 to 2022.
-We find that $Dis$ reduces TFP, especially in state-owned, non-capital-intensive, and low-pollution firms.
-Mechanism analysis shows that green innovation strengthens the dampening effect of $Dis$ on TFP, and that $Dis$ lowers corporate TFP by increasing financing constraints.
+We find that $Dis$ reduces TFP, especially in state-owned, non-capital-intensive, low-pollution and high-tech firms, green innovation strengthens the dampening effect of $Dis$ on TFP, and that $Dis$ lowers corporate TFP by increasing financing constraints and weakening human capital.
 Furthermore, XGBoost regression demonstrates that $Dis$ plays a significant role in predicting TFP, with SHAP showing that the dampening effect of ESG rating disagreement on TFP is still pronounced in firms with large $Dis$ values.
 JEL Codes: G32, G24, C54.
-Keywords: ESG rating disagreement; Total factor productivity; Green innovation; Financing constraints; Machine learning
+Keywords: ESG rating disagreement; Total factor productivity; Green innovation; Financing constraints; Human Capital; Machine learning
 1 Introduction
@@
 Conversely, significant ESG rating disagreements can intensify information asymmetry, creating a noise effect that leads to investor misjudgments(Berg et al., 2022).
-Thereby, increasing financing constraints and ultimately resulting in a decline in total factor productivity.
+Thereby, increasing financing constraints and ultimately resulting in a decline in TFP.
 Simultaneously, based on stakeholder theory, companies may take certain remedial measures in response to ESG rating disagreements to balance the interests of various stakeholders.
-For example, companies might intensify green innovation under the stimulus of ESG rating disagreements(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); However, as this initiative cannot cater to all stakeholders under noise effect caused by ESG rating disagreement, it may lead some groups to perceive "greenwashing" behaviors, which ultimately undermines the company’s financing efforts and results in a decline in total factor productivity(Peng and Xie, 2024).
+Companies might intensify green innovation under the stimulus of ESG rating disagreements(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); However, as this initiative cannot cater to all stakeholders under noise effect caused by ESG rating disagreement, it may lead some groups to perceive "greenwashing" behaviors, which ultimately undermines the company’s financing efforts and results in a decline in TFP(Peng and Xie, 2024).
+Meanwhile, the divergence in ESG ratings may also drive companies to focus on improving their ESG recognition, potentially causing an imbalance between the interests of external investors and internal employees.
+This imbalance could lead to a human capital outflow, ultimately affecting the development of TFP.
 In summary, we propose the research hypothesis H1: ESG rating disagreement reduces corporate TFP.
@@
 3.1 Sample Selection
-We utilizes financial data of A-share listed companies in China from 2015 to 2022111Before 2015, the only more well-known local ESG rating agencies in China were Huazheng, alongside ESG ratings from four agencies: Huazheng, Wind, SynTao Green Finance, and MioTech.
+We utilize financial data of A-share listed companies in China from 2015 to 2022111Prior 2015, the widely recognized local ESG rating agency in China was Huazheng., alongside ESG ratings from four agencies: Huazheng, Wind, SynTao Green Finance, and Susallwave.
 The data processing procedures follow these steps: (1) excluding companies in the financial and real estate industries; (2) excluding ST, *ST, and delisted companies; (3) excluding companies with only one ESG rating in a given year; (4) excluding companies with missing key indicators; and (5) applying Winsorization at the 1% level on both tails.
-All data is sourced from the CSMAR database, the CNRDS database, and the four rating agencies, resulting in 13417 samples in baseline regression.
+All data is sourced from the CSMAR database, the CNRDS database, and the four rating agencies, resulting in 3656 firms and 13417 samples in baseline regression.
 3.2 Model Specification
-To test the hypotheses presented earlier,we conduct the following baseline regression model [^2] Considering the time required for information to be transmitted and for the market to react, TFP is lagged by one period:
+To test the hypotheses presented earlier,we conduct the following baseline regression model [^2] Considering the time required for information to be transmitted and for the market to react and mitigate potential reverse causation problems, TFP is lagged by one period.:
 $TFP_{i,t+1}=\alpha_{0}+\alpha_{1}Dis_{i,t}+\alpha_{2}Control_{i,t}+Year_{t}+Id_{i}+\epsilon_{i,t}$ (1)
@@
 Domestic ESG rating agencies have advantages in terms of time and spatial proximity that foreign agencies cannot match, making them more likely to influence the behavior of domestic firms.
-Therefore, this study uses ESG rating data from four Chinese agencies: Huazheng, Wind, SynTao Green Finance, and MioTech.
+Therefore, this study uses ESG rating data from four Chinese agencies: Huazheng, Wind, SynTao Green Finance, and Susallwave.
 Following Avramov et al.
@@
 $FC$ CSAMR WW and KZ indices, the larger they are the more constrained the financing is
+$HC$ CSAMR Natural logarithm of number of graduate students and above and natural logarithm of number of undergraduate students and above
 4 Empirical Analysis
-Table 2: Descriptive Statistics of Main Variables
-Variables Count Mean SD Min 25% 50% 75% Max
-$TFP_{LP}$ 13417 8.51 1.04 6.38 7.79 8.40 9.14 11.28
-$TFP_{OP}$ 13417 6.83 0.85 5.14 6.24 6.72 7.33 9.17
-$TFP_{GMM}$ 13417 5.72 0.80 4.09 5.18 5.63 6.18 8.12
-$Dis$ 13417 0.26 0.19 0.01 0.11 0.24 0.34 0.83
-$TobinQ$ 13417 2.03 1.32 0.83 1.22 1.61 2.31 8.34
-$Top10$ 13417 56.77 15.15 23.59 45.91 56.95 67.87 90.89
-$ListAge$ 13417 2.34 0.68 1.10 1.79 2.40 3.00 3.37
-$Lev$ 13417 0.43 0.19 0.06 0.28 0.43 0.57 0.88
-$Size$ 13417 22.52 1.34 20.09 21.54 22.32 23.31 26.41
-$Avg$ 13417 4.79 0.80 3.07 4.25 4.69 5.25 7.17
-$Der$ 13417 1.05 1.07 0.07 0.39 0.74 1.31 6.75
-$EI$ 13417 1.30 1.49 0.00 0.00 1.10 2.30 7.74
-$WW$ 10059 -1.03 0.08 -1.24 -1.08 -1.03 -0.98 -0.86
-$KZ$ 13417 1.17 1.94 -4.35 0.02 1.36 2.49 6.22
-  • Note: The data in the table are descriptive statistics for deleting missing values after lagging one period of TFP.
 4.1 Baseline Regression
 Table 2 reports the results of the baseline regression, with clustering at the individual and time levels.
-The results indicate that ESG rating disagreement reduces corporate total factor productivity, providing preliminary support for the validity of Hypothesis H1.
+The results indicate that ESG rating disagreement reduces corporate TFP, providing preliminary support for the validity of Hypothesis H1.
 Table 2: Baseline Regression Results
@@
   • Note: t-statistics are in parentheses. * p<0.1, ** p<0.05, *** p<0.01. The same applies below.
-4.2 Robustness Tests
-To ensure the robustness of the baseline regression results, the following robustness tests were conducted.
-First, to verify whether Hypothesis H1 holds when only leading $Dis$ by one period, we lead the it by one period(see table8).
-Second, since the total factor productivity of firms in this study is influenced by regional productivity levels, to mitigate endogeneity issues arising from omitted variable bias, we include the provincial-level New quality productivity measured by Lu et al.
-(2024) as a control variable (see table9).
-Third, following Yang et al.
-(2022), we calculate the industry average of ESG rating disagreements by year and industry, and use the cube of the difference between the industry average and the ESG rating disagreements as an instrumental variable.
-A two-stage least squares (2SLS) estimation is then conducted to further mitigate potential endogeneity issues (see table10).
-Fourth, we denote disagreements under 20% as non-existent and the remainder as present.
-Propensity scores are computed via logistic regression, followed by 1:1 nearest-neighbor matching and subsequent regression (see table11).
 4.2 Mechanism Analysis
@@
 Conversely, when there are ESG rating discrepancies, it strengthens the dampening effect of ESG rating disagreement on TFP.
-Dose this contradicts exiting researches that show firms tend to increase their green innovation efforts when ESG rating disagreement is present(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b).
+The result seem contradict exiting researches, which argue ESG rating disagreement can drive firms to enhance green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b).
 This observation is very likely caused by a remedial action taken by firms in response to ESG rating disagreement, as Hou and Xie (2024) mentioned, such ESG rating disagreement is negatively correlated with the quality of corporate green innovation.
@@
 Figure 1: Marginal Effect
+To further investigate whether the moderating effect varies across levels of ESG rating disagreement, we conducted regressions separately for the upper and lower 50% quartiles of ESG rating disagreement, and the tabel 4 show that the moderating effect is significant only when the ESG rating disagreement is high, further supporting the notion that substantial ESG rating disagreement may lead to stakeholder distrust in firms’ green innovation.
+Table 4: Heterogeneity of Green Innovation Mechanism
+Variables TFP
+High $Dis$ Low $Dis$
+$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
+$Dis$ 0.0527 0.0616 0.0501 -0.2268*** -0.1532** -0.1554**
+(1.2360) (1.5631) (1.2368) (-2.8516) (-2.0471) (-1.9903)
+$EI$ 0.0364*** 0.0280*** 0.0230** 0.0092 0.0043 0.0041
+(3.1495) (2.6193) (2.0914) (1.4601) (0.7258) (0.6628)
+$Dis*EI$0.0522** -0.0461** -0.0381** 0.0433 0.0280 0.0260
+(-2.1556) (-2.0558) (-1.6525) (1.2832) (0.8798) (0.7855)
+$TobinQ$0.0009 -0.0012 0.0020 -0.0025 -0.0020 -0.0013
+(-0.1395) (0.1681) (0.3447) (-0.4475) (-0.3853) (-0.2318)
+$Top10$0.0008 -0.0012 -0.0005 -0.0001 0.0004 0.0008
+(-0.8389) (-1.3029) (-0.5405) (-0.1141) (0.4905) (0.8934)
+$ListAge$ 0.1137** 0.0833* 0.0257 0.0609 0.0529 -0.0332
+(2.2567) (1.7886) (0.5359) (1.4289) (1.3186) (-0.7922)
+$Lev$0.1140 -0.0997 -0.0976 -0.2786*** -0.2703*** -0.2712***
+(0.1716) (-1.2930) (-1.2306) (-3.2945) (-3.3962) (-3.2670)
+$Size$ 0.3607*** 0.2525*** 0.2173*** 0.3523*** 0.2355*** 0.2123***
+(16.816) (12.732) (10.655) (17.502) (12.434) (10.746)
+$Avg$ 0.2126** 0.2424*** 0.2447*** 0.2301*** 0.2706*** 0.2651***
+(13.198) (16.277) (15.983) (13.969) (17.451) (16.389)
+$Der$ 0.0224** -$0*$0226** 0.0046 0.0052 0.0036
+(-2.2377) (-2.2735) (-2.3751) (0.3638) (0.4404) (0.2910)
+Controls YES YES YES YES YES YES
+Year YES YES YES YES YES YES
+Id YES YES YES YES YES YES
+$N$ 6774 6774 6774 6643 6643 6643
+$R^{2}$ 0.1578 0.1434 0.1229 0.1717 0.1533 0.1310
 4.2.2 The Role of Financing Constraints
@@
 Table 5: Financing Constraints Mechanism
-Variables WW
+Variables Financing Constraints
 $WW$ $WW$ $KZ$ $KZ$ $WW$ $WW$ $KZ$ $KZ$
@@
 $TobinQ$0.0022*** -0.0006 0.2724*** 0.0237
-(-3.9269) -0.9530 (19.528) (1.4433)
+(-3.9269) (-0.9530) (19.528) (1.4433)
 $Top10$0.0007*** -0.0006*** 0.0011 -0.0061**
-(-8.8614) -6.9840 (0.4783) (-2.3496)
+(-8.8614) (-6.9840) (0.4783) (-2.3496)
 $ListAge$ 0.0204*** 0.0143*** 0.5530*** 1.2048***
@@
 (3.4646) (1.2994) (0.6820) (-0.0931)
-Controls NO NO NO YES YES YES YES YES
+Controls NO NO NO NO YES YES YES YES
 Year YES YES YES YES YES YES YES YES
@@
   • Note: Similar to the robustness tests, $Dis$ corresponds to regressions on ESG rating disagreement in year t and financing constraints in year t+1, while $Disb$ corresponds to regressions on ESG rating disagreement in year t-1 and financing constraints in year t.
+4.2.3 The Role of Human Capital
+Human capital is widely acknowledged as a positive contributor to TFP(Miller and Upadhyay, 2000; Xu et al., 2008; Huang et al., 2024).
+Employees with a bachelor’s or master’s degree typically reflects higher levels of knowledge, skills, and expertise among workers, enabling them to perform task more effciently, reduce errors, and enhance productivity.
+In this subsection, we delve deeper into how ESG rating disagreement interacts with human capital and its potential impact on a company’s TFP, with a specific focus on employees holding bachelor’s or master’s degrees.
+Building on Jiang (2022)’s research on the transmission mechanism, and following a similar approach to verifying the financing constraint mechanism to avoid endogeneity issues in mediating variables, we perform regressions of $Dis$ on $Master$ and $Dis$ on $Undergrad$.
+Table 6 reports the regression results, showing that Dis is detrimental to the development of firms’ human capital, which serves as a crucial driver of TFP.
+Based on this evidence, we argue that Dis reduces TFP by impairing human capital.
+Table 6: Human Capital Mechanism
+Variables Human Capital
+$Master$ $Master$ $Undergrad$ $Undergrad$ $Master$ $Master$ $Undergrad$ $Undergrad$
+$Dis$0.0595*** -0.0526*** -0.0345** -0.0316**
+(-2.8695) (-3.1495) (-1.7625) (-2.0461)
+$Disb$0.0598*** -0.0528** -0.0322* -0.0248*
+(-2.8867) (-3.1608) (-1.8268) (-1.9302)
+$TobinQ$ 0.0197*** -0.0067 0.2724*** -0.0036
+(5.0228) (-1.4155) (19.528) (-0.9603)
+$Top10$ 0.0029*** 0.0008 0.0011 0.0003
+(4.4620) (1.0079) (0.4783) (0.4613)
+$ListAge$0.0010 0.1365*** 0.5530*** 0.1578***
+(-0.0413) (3.7156) (6.5333) (5.5267)
+$Lev$ 0.0536 -0.0620 6.7446*** 0.0894*
+(0.8913) (-0.9460) (32.842) (1.7679)
+$Size$ 0.6722*** 0.4376*** -0.4512*** 0.4410***
+(49.844) (27.508) (-9.5683) (35.103)
+$Avg$0.1597*** -0.0907*** -0.5288*** -0.0739***
+(-14.054) (-7.0579) (-13.531) (-7.5015)
+$Der$0.0256*** -0.0128 0.0191 -0.0184***
+(-3.1008) (-1.5259) (0.6820) (-2.8025)
+Controls NO NO NO NO YES YES YES YES
+Year YES YES YES YES YES YES YES YES
+Id YES YES YES YES YES YES YES YES
+$N$ 10461 10461 12595 12595 10461 10461 12595 12595
+$R^{2}$ 0.0011 0.0011 0.0010 0.0011 0.2790 0.1145 0.4112 0.1511
+  • Note: Similar to the robustness tests, $Dis$ corresponds to regressions on ESG rating disagreement in year t and human capital in year t+1, while $Disb$ corresponds to regressions on ESG rating disagreement in year t-1 and human capital in year t.
 4.3 Heterogeneity Analysis
@@
 Table 7 reports the results of the heterogeneity analysis.
-The results indicate that the effect is more significant in state-owned firms, non-capital-intensive firms, and low-pollution firms.
-The above heterogeneity may exist because, first, there is a greater willingness to engage in ESG construction among non-state-owned firms due to the fact that they tend to focus more on ESG performance to attract more investment; second, usually capital-intensive firms are less affected by market forces due to their own low variable costs and high fixed costs; third, in the case of highly polluting firms that are susceptible to a variety of environmental policies and regulations in their financing, such firms are less likely to invest in ESG performance, and therefore the impact of divergent ESG ratings is less significant.
+The results indicate that the effect is more significant in state-owned firms, non-capital-intensive firms, low-pollution firms and high-tech firms.
+This heterogeneity may arise for the following reasons: first, non-state-owned firms are generally more willing to engage in ESG initiatives as they tend to focus more on ESG performance to attract investment; second, capital-intensive firms are usually less influenced by market forces due to their low variable costs and high fixed costs, which buffer them from the pressures of ESG-related changes; third, highly polluting firms, which are often subject to stringent environmental policies and regulations in their financing, are less likely to invest in ESG performance, making the impact of divergent ESG ratings less pronounced.
+And, finally, as previously analyzed, ESG rating disagreements can lead to the loss of highly qualified employees, on whom high-tech firms are particularly dependent.
+Consequently, the negative impact of disagreements on TFP is more significant for high-tech firms.
 Table 7: Heterogeneity Analysis
 Variables TFP
-State-Owned Non-State-Owned Capital-Intensive Non-Capital-Intensive High-Pollution Low-Pollution
+State-Owned Non-State-Owned Capital-Intensive Non-Capital-Intensive High-Pollution Low-Pollution High-Tech Non-High-Tech
 $TFP_{LP}$
-$Dis$0.0622** -$0$0322 -0.0462*** -0.0239 -0.0481***
-(-2.0516) (-1.8294) (-1.1009) (-2.8884) (-1.0401) (-2.7959)
-Chow Test 500.19*** 289.12*** 544.57***
-$R^{2}$ 0.1573 0.1886 0.1741 0.1817 0.2010 0.1754
+$Dis$0.0622** -$0$0322 -0.0462*** -0.0239 -0.0481*** -0.0425** -0.0043
+(-2.0516) (-1.8294) (-1.1009) (-2.8884) (-1.0401) (-2.7959) (-2.2163) (-0.1653)
+Chow Test 500.19*** 289.12*** 544.57*** 756.47***
+$R^{2}$ 0.1573 0.1886 0.1741 0.1817 0.2010 0.1754 0.1751 0.1570
 $TFP_{OP}$
-$Dis$0.0553* -0.0268 -0.0097 -0.0458*** -0.0070 -0.0493**
-(-1.8608) (-1.5208) (-0.3229) (-2.6338) (-0.3149) (-2.4735)
-Chow Test 529.58*** 269.31*** 534.66***
-$R^{2}$ 0.1759 0.1593 0.1672 0.1663 0.1930 0.1595
+$Dis$0.0553* -0.0268 -0.0097 -0.0458*** -0.0070 -0.0493** -0.0365** -0.0048
+(-1.8608) (-1.5208) (-0.3229) (-2.6338) (-0.3149) (-2.4735) (-2.2763) (-0.1639)
+Chow Test 529.58*** 269.31*** 534.66*** 760.50***
+$R^{2}$ 0.1759 0.1593 0.1672 0.1663 0.1930 0.1595 0.1561 0.1526
 $TFP_{GMM}$
-$Dis$0.0618** -$0$0278 -0.0471*** -0.0134 -0.0519**
-(-2.1381) (-1.6745) (-0.9906) (-2.6889) (-0.6185) (-2.5484)
-Chow Test 498.97*** 248.15*** 475.42***
-$R^{2}$ 0.1607 0.1254 0.1504 0.1377 0.1691 0.1321
-Controls YES YES YES YES YES YES
-Year YES YES YES YES YES YES
-Id YES YES YES YES YES YES
-Double Clustering YES YES YES YES YES YES
-$N$ 4416 9001 2342 11075 3900 9517
+$Dis$0.0618** -$0$0278 -0.0471*** -0.0134 -0.0519** -0.0413** -0.0122
+(-2.1381) (-1.6745) (-0.9906) (-2.6889) (-0.6185) (-2.5484) (-2.4165) (-0.4205)
+Chow Test 498.97*** 248.15*** 475.42*** 698.54***
+$R^{2}$ 0.1607 0.1254 0.1504 0.1377 0.1691 0.1321 0.1256 0.1315
+Controls YES YES YES YES YES YES YES YES
+Year YES YES YES YES YES YES YES YES
+Id YES YES YES YES YES YES YES YES
+Double Clustering YES YES YES YES YES YES YES YES
+$N$ 4416 9001 2342 11075 3900 9517 8454 4963
 5 Further Discussion
@@
 The variables include the ESG rating disagreement ($Dis$ $Disb$), all control variables, and institutional ownership percentage ($Inst$).
-We use 80% of the dataset for training and 20% for testing and conducted three rounds of training: the first round included ESG rating disagreement, the second round added ESG rating disagreement from the previous period, and the third round did not include ESG rating disagreement.
-Table 7 reports the evaluation metrics for the three rounds of training.
-It can be observed that using $R^{2}$ as the evaluation metric, the model’s predictive performance improved by 4.3% and 6.23% respectively compared to the model without ESG rating disagreement, indicating that ESG rating disagreement aids in the short-term prediction of $TFP_{LP}$.
+We use 80% of the dataset for training and 20% for testing, conducting three rounds of training.
+The first round includes ESG rating disagreement (Dis), the second round incorporates ESG rating disagreement from the previous period, and the third round excludes ESG rating disagreement altogether.
+Table 8 presents the evaluation metrics for these three rounds of training.
+First, compared to the linear model (Set4), XGBoost demonstrates significantly better performance across all metrics, highlighting the machine learning model’s ability to capture nonlinear relationships.
+Second, when comparing XGBoost without incorporating Dis (Set3), the model’s performance improves by 4.3% (Set1) and 6.23% (Set2), respectively, using $R^{2}$ as the evaluation criterion.
+This improvement underscores the predictive value of ESG ratings in modeling TFP.
 Table 8: Performance Metrics for Training and Test Sets
-Metric Training Set 1 Training Set 2 Training Set 3 Test Set 1 Test Set 2 Test Set 3
-MSE 0.2609 0.2274 0.2818 0.3492 0.3300 0.3835
-RMSE 0.5108 0.4769 0.5309 0.5909 0.5744 00.6192
-MAE 0.3995 0.3729 0.4145 0.4589 0.4523 0.4818
-MAPE 4.7786% 4.4569% 4.9457% 5.4649% 5.3644% 5.7293%
-R2 0.7685 0.7969 0.7500 0.7045 0.7176 0.6755
-To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in the predictive task, SHAP was used to interpret the XGBoost model.
-SHAP values, rooted in game theory, quantify each feature’s contribution to the model output by treating features as "players" in a game.
-The Shapley value $\phi_{i}$ for a feature $i$ is calculated as:
+Metric Training Set 1 Training Set 2 Training Set 3 Training Set 4 Test Set 1 Test Set 2 Test Set 3 Test Set 4
+MSE 0.2609 0.2274 0.2818 0.4714 0.3492 0.3300 0.3835 0.4728
+RMSE 0.5108 0.4769 0.5309 0.6866 0.5909 0.5744 0.6192 0.6876
+MAE 0.3995 0.3729 0.4145 0.5383 0.4589 0.4523 0.4818 0.5415
+MAPE 4.7786% 4.4569% 4.9457% 6.4047% 5.4649% 5.3644% 5.7293% 6.4399%
+R2 0.7685 0.7969 0.7500 0.5794 0.7045 0.7176 0.6755 0.5955
+  • Note: Set1 is the XGBoost performance of $Dis$ in the same period as the TFP, Set2 is the XGBoost performance of $Dis$ in the preceding period, Set3 is the XGBoost performance excluding $Dis$, and Set4 is the model performance of the linear model (no fixed effects) and $Dis$ in the preceding period.
+To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in our predictive task, we applied SHAP to interpret the results of the XGBoost model.
+SHAP values are grounded in cooperative game theory and provide a way to understand each feature’s contribution to the model output by treating features as "players" in a game.
+The contribution, or Shapley value $\phi_{i}$, for a particular feature $i$ is computed by averaging its marginal contributions across all possible combinations of features.
+Mathematically, the Shapley value $\phi_{i}$ for feature $i$ is defined as:
 $\phi_{i}=\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!\cdot(|N|-|S|-1)!}{|N|!}\left[f(S\cup\{i\})-f(S)\right]$ (3)
-where $S$ is a subset excluding $i$, $f(S)$ is the model output for $S$, and $f(S\cup\{i\})$ is the output after adding $i$ to $S$.
+In this equation:
+  • $N$ represents the set of all features, while $S$ is a subset of $N$ that does not include feature $i$.
+  • $|S|$ denotes the size of subset $S$.
+  • $f(S)$ is the model output based on the subset of features $S$.
+  • $f(S\cup\{i\})$ is the output when feature $i$ is added to $S$.
+The term $\frac{|S|!\cdot(|N|-|S|-1)!}{|N|!}$ in the equation represents a weighting factor that accounts for all possible orders in which features can be added to the model.
+By calculating $\phi_{i}$ across all possible subsets $S$, SHAP captures the average marginal effect of adding feature $i$ to different subsets of features, thereby quantifying the feature’s contribution to the model’s output.
 In this study, based on the previous training results, we chose the model with one-period lagged ESG rating disagreement for our SHAP analysis, and the SHAP values were calculated for the test set.
-Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting total factor productivity.
-Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s total factor productivity, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task.
+Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting TFP [^3] The importance of each feature is calculated as the percentage of the average of the absolute SHAP values from each to the mean of the absolute SHAP values of all features; in short, a higher percentage means that the feature is more important among all features..
+Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s TFP, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task [^4] The beeswarm plot shows the effect of different values of feature on TFP, the color of the dots in the plot shifts toward red for larger eigenvalues and toward blue for smaller eigenvalues and points positioned on the horizontal axis with values less than 0 indicate a negative effect on the dependent variable in the model, while points with values greater than 0 signify a positive effect..
 The reason for this phenomenon is that the investment market is more tolerant of small ESG rating divergence, and listed companies will not face significant financing pressure due to small divergence.
@@
 Figure 3: SHAP Beeswarm Plot
-6 Conclusion
-ESG rating disagreement poses challenges to the enhancement of corporate TFP.
-This study utilizes financial data from China’s A-share listed companies and ESG rating information from domestic ESG rating agencies between 2015 and 2022 to explore the relationship between ESG rating disagreements and firms’ TFP.
-The findings robustly indicate that ESG rating disagreements have a negative impact on firms’ TFP.
-The mechanism analysis reveals that green innovation instead strengthens the dampening effect of ESG rating disagreement on TFP; moreover, ESG rating disagreement strengthens firms’ financing constraints, thereby hindering the improvement of TFP.
-Heterogeneity analysis further indicates that the negative impact of ESG rating disagreements on TFP is more pronounced in state-owned enterprises, non-capital-intensive firms, and companies with lower levels of pollution.
-Additionally, an important finding from the analysis using machine learning methods suggests that ESG rating disagreements serve as a significant indicator for short-term predicting firms’ TFP, providing new insights for companies to engage in ESG risk management to enhance TFP.
+6 Conclusion and Implication
+ESG rating disagreements present significant challenges to the improvement of corporate TFP, underscoring the urgent need for firms to address this issue.
+Mechanism analysis reveals that green innovation can exacerbate the negative effects of ESG rating disagreements on TFP.
+Additionally, these disagreements hinder TFP growth by intensifying financing constraints and weakening human capital.
+To mitigate these impacts, companies should focus on improving the quality of ESG information disclosure, carefully navigate the trade-offs between green innovation, financing needs, and strive to balance the interests of diverse stakeholders.
+Regulatory authorities, in turn, should strengthen the scrutiny of green innovation patents to ensure their quality and prevent superficial compliance efforts.
+Heterogeneity analysis demonstrates that the negative effects of ESG rating disagreements on TFP are more pronounced in state-owned enterprises, non-capital-intensive firms, low-pollution industries, and high-tech enterprises.
+These entities should pay closer attention to ESG rating discrepancies, while regulators should provide targeted support to low-pollution and high-tech firms to foster sustainable development and technological innovation.
+Moreover, ESG rating disagreements serve as a critical predictor of TFP, with SHAP values highlighting their substantial and ongoing negative impact.
+This indicator can be leveraged by listed companies and regulatory agencies as a monitoring and early warning tool to identify potential risks to TFP and implement proactive measures.
 Appendix A Appendix
-A.1 Robustness Tests
+A.1 Descriptive Statistics
+Table 9: Descriptive Statistics of Main Variables
+Variables Count Mean SD Min 25% 50% 75% Max
+$TFP_{LP}$ 13417 8.51 1.04 6.38 7.79 8.40 9.14 11.28
+$TFP_{OP}$ 13417 6.83 0.85 5.14 6.24 6.72 7.33 9.17
+$TFP_{GMM}$ 13417 5.72 0.80 4.09 5.18 5.63 6.18 8.12
+$Dis$ 13417 0.26 0.19 0.01 0.11 0.24 0.34 0.83
+$TobinQ$ 13417 2.03 1.32 0.83 1.22 1.61 2.31 8.34
+$Top10$ 13417 56.77 15.15 23.59 45.91 56.95 67.87 90.89
+$ListAge$ 13417 2.34 0.68 1.10 1.79 2.40 3.00 3.37
+$Lev$ 13417 0.43 0.19 0.06 0.28 0.43 0.57 0.88
+$Size$ 13417 22.52 1.34 20.09 21.54 22.32 23.31 26.41
+$Avg$ 13417 4.79 0.80 3.07 4.25 4.69 5.25 7.17
+$Der$ 13417 1.05 1.07 0.07 0.39 0.74 1.31 6.75
+$EI$ 13417 1.30 1.49 0.00 0.00 1.10 2.30 7.74
+$WW$ 10059 -1.03 0.08 -1.24 -1.08 -1.03 -0.98 -0.86
+$KZ$ 13417 1.17 1.94 -4.35 0.02 1.36 2.49 6.22
+$Master$ 10461 4.31 1.55 1.10 3.18 4.20 5.31 8.42
+$Undergrad$ 12595 6.31 1.35 3.56 5.35 6.17 7.15 10.05
+  • Note: The data in the table are descriptive statistics for deleting missing values after lagging one period of TFP.
+A.2 Correlation Coefficient
+Table 10: Correlation Matrix
+$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $Dis$ $TobinQ$ $Top10$ $ListAge$ $Lev$ $Size$ $Avg$ $Der$ $N.TFP_{province}$ $IV$
+$TFP_{LP}$ 1.000 0.951 0.918 -0.066 -0.200 0.127 0.329 0.444 0.801 0.648 0.344 -0.037 0.082
+$TFP_{OP}$ 0.951 1.000 0.976 -0.054 -0.183 0.104 0.303 0.405 0.707 0.841 0.325 -0.036 0.067
+$TFP_{GMM}$ 0.918 0.976 1.000 -0.041 -0.132 0.080 0.244 0.356 0.583 0.823 0.295 -0.005 0.059
+$Dis$0.066 -0.054 -0.041 1.000 0.014 -0.078 0.025 0.039 -0.071 -0.024 0.052 -0.015 -0.974
+$TobinQ$0.200 -0.183 -0.132 0.014 1.000 -0.036 -0.094 -0.249 -0.283 -0.116 -0.173 0.012 -0.008
+$Top10$ 0.127 0.104 0.080 -0.078 -0.036 1.000 -0.364 -0.066 0.147 0.044 -0.052 0.038 0.087
+$ListAge$ 0.329 0.303 0.244 0.025 -0.094 -0.364 1.000 0.292 0.424 0.204 0.240 -0.184 -0.024
+$Lev$ 0.444 0.405 0.356 0.039 -0.249 -0.066 0.292 1.000 0.458 0.256 0.843 -0.044 -0.038
+$Size$ 0.801 0.707 0.583 -0.071 -0.283 0.147 0.424 0.458 1.000 0.413 0.359 -0.104 0.082
+$Avg$ 0.648 0.841 0.823 -0.024 -0.116 0.044 0.204 0.256 0.413 1.000 0.222 -0.041 0.027
+$Der$ 0.344 0.325 0.295 0.052 -0.173 -0.052 0.240 0.843 0.359 0.222 1.000 -0.047 -0.046
+$N.TFP_{province}$0.037 -0.036 -0.005 -0.015 0.012 0.038 -0.184 -0.044 -0.104 -0.041 -0.047 1.000 0.014
+$IV$ 0.082 0.067 0.059 -0.974 -0.008 0.087 -0.024 -0.038 0.082 0.027 -0.046 0.014 1.000
+A.3 Robustness Tests and Endogeneity Treatment
+To ensure the robustness of the baseline regression results, the following robustness tests were conducted.
+First, to verify whether Hypothesis H1 holds when only leading $Dis$ by one period, we lead the it by one period(see table 11).
+Second, since the TFP of firms in this study is influenced by regional productivity levels, to mitigate endogeneity issues arising from omitted variable bias, we include the provincial-level New quality productivity measured by Lu et al.
+(2024) as a control variable (see table 12).
+Third, to further mitigate everse causality endogeneity issues, following Yang et al.
+(2022), we calculate the industry average of ESG rating disagreements by year and industry, and use the cube of the difference between the industry average and the ESG rating disagreements as an instrumental variable.
+A two-stage least squares (2SLS) estimation is conducted, and the Kleibergen-Paap rk LM statistics verify the verify the validity of instrumental variables(see table 13).
+Fourth, to address sample selection bias, we denote disagreements under 20% as non-existent and the remainder as present.
+Propensity scores are computed via logistic regression, followed by 1:1 nearest-neighbor matching and subsequent regression (see table 14).
 Table 11: Robustness Tests (I)
@@
 $Dis$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
-$IV(\hat{Dis})$0.3336*** -0.0479*** -0.0408** -0.0424**
-(-602.14) (-3.0748) (-2.3966) (-2.3533)
-$TobinQ$0.0007 -0.0027 -0.0013 -0.0004
-(-0.5971) (-0.4037) (-0.2634) (-0.0756)
-$Top10$ 3.476e-05 -0.0010 -0.0009 -0.0005
-(0.3084) (-0.9808) (-0.9708) (-0.5052)
-$ListAge$0.0056 0.0910** 0.0799** 0.0061
-(-0.9005) (2.4981) (2.3788) (0.1734)
-$Lev$0.0029 -0.1554 -0.1407 -0.1333
-(-0.3495) (-1.3058) (-1.2618) (-1.1844)
-$Size$0.0045** 0.3760*** 0.2594*** 0.2260***
-(-2.6017) (14.301) (13.453) (10.029)
-$Avg$0.0017 0.2278*** 0.2680** 0.2660**
-(-0.7314) (2.6949) (2.7600) (2.6811)
-$Der$ 0.0037** -$0$0193 -0.0216
-(2.8094) (-1.9412) (-1.7701) (-1.9119)
-$N.TFP_{province}$ 0.0100 0.0999 0.0851 0.0408
-(0.5583) (0.6350) (0.5902) (0.2565)
+$IV(\hat{Dis})$0.3302*** -0.0476*** -0.0409** -0.0425**
+(-328.74) (-3.0292) (-2.3843) (-2.3456)
+$TobinQ$ 0.0002 -0.0027 -0.0013 -0.0003
+(0.1447) (-0.3986) (-0.2574) (-0.0685)
+$Top10$ 0.0001 -0.0009 -0.0009 -0.0005
+(1.1390) (-0.9769) (-0.9679) (-0.5016)
+$ListAge$0.0118 0.0906** 0.0797** 0.0058
+(-1.2204) (2.4886) (2.3696) (0.1655)
+$Lev$0.0061 -0.1555 -0.1408 -0.1335
+(-0.7749) (-1.3069) (-1.2627) (-1.1853)
+$Size$0.0044** 0.3761*** 0.2594*** 0.2260***
+(-2.0583) (14.306) (13.451) (10.030)
+$Avg$0.0015 0.2277*** 0.2680** 0.2660**
+(-0.5293) (2.6949) (2.7600) (2.6812)
+$Der$ 0.0045** -$0$0193 -0.0216
+(3.1346) (-1.9398) (-1.7686) (-1.9099)
+$N.TFP_{province}$0.0030 0.0994 0.0847 0.0404
+(-0.2042) (0.6320) (0.5877) (0.2540)
 Controls YES YES YES YES
@@
 Double Clustering YES YES YES YES
-$F$ 2.983e+04 227.22 208.52 170.30
-$N$ 16907 12892 12892 12892
-R2 0.9542 0.1794 0.1671 0.1408
-  • Note: The second stage of 2SLS regression follows the baseline regression with a one-period lag.
+$F$ 2.049e+04 227.20 208.52 170.30
+$N$ 12892 12892 12892 12892
+Kleibergen-Paap rk LM 622.33****
+R2 0.9517 0.1794 0.1671 0.1408
+  • Note: TFP is t+1 period, just like the baseline regression setup.
 Table 14: Robustness Tests (IV)
@@
 R2 0.0028 0.0027 0.0024 0.1739 0.1628 0.1455
-A.2 Optuna Parameter Range & Correlation Coefficient
-Table 12: Correlation Matrix
-$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $Dis$ $TobinQ$ $Top10$ $ListAge$ $Lev$ $Size$ $Avg$ $Der$ $N.TFP_{province}$ $IV$
-$TFP_{LP}$ 1.000 0.951 0.918 -0.066 -0.200 0.127 0.329 0.444 0.801 0.648 0.344 -0.037 0.082
-$TFP_{OP}$ 0.951 1.000 0.976 -0.054 -0.183 0.104 0.303 0.405 0.707 0.841 0.325 -0.036 0.067
-$TFP_{GMM}$ 0.918 0.976 1.000 -0.041 -0.132 0.080 0.244 0.356 0.583 0.823 0.295 -0.005 0.059
-$Dis$0.066 -0.054 -0.041 1.000 0.014 -0.078 0.025 0.039 -0.071 -0.024 0.052 -0.015 -0.974
-$TobinQ$0.200 -0.183 -0.132 0.014 1.000 -0.036 -0.094 -0.249 -0.283 -0.116 -0.173 0.012 -0.008
-$Top10$ 0.127 0.104 0.080 -0.078 -0.036 1.000 -0.364 -0.066 0.147 0.044 -0.052 0.038 0.087
-$ListAge$ 0.329 0.303 0.244 0.025 -0.094 -0.364 1.000 0.292 0.424 0.204 0.240 -0.184 -0.024
-$Lev$ 0.444 0.405 0.356 0.039 -0.249 -0.066 0.292 1.000 0.458 0.256 0.843 -0.044 -0.038
-$Size$ 0.801 0.707 0.583 -0.071 -0.283 0.147 0.424 0.458 1.000 0.413 0.359 -0.104 0.082
-$Avg$ 0.648 0.841 0.823 -0.024 -0.116 0.044 0.204 0.256 0.413 1.000 0.222 -0.041 0.027
-$Der$ 0.344 0.325 0.295 0.052 -0.173 -0.052 0.240 0.843 0.359 0.222 1.000 -0.047 -0.046
-$N.TFP_{province}$0.037 -0.036 -0.005 -0.015 0.012 0.038 -0.184 -0.044 -0.104 -0.041 -0.047 1.000 0.014
-$IV$ 0.082 0.067 0.059 -0.974 -0.008 0.087 -0.024 -0.038 0.082 0.027 -0.046 0.014 1.000
+A.4 Optuna Parameter Range
 Table 15: Hyperparameters for XGBoost Model

### v5

@@
 1 Introduction
-Since 1987, the United Nations has reinforced its commitment to sustainable development, and introduced the concept of ESG (Environmental, Social, and Governance) by then-UN Secretary-General Kofi Annan in 2004.
-Since then, ESG principles have been increasingly integrated into global investment strategies and corporate assessments, with their specific meanings evolving over time.
-In recent years, China is prominently advocating the philosophy that "green mountains and clear waters are as valuable as mountains of gold and silver", and promoting sustainable development both domestically and internationally.
-On December 27, 2023, the Guidelines for the Comprehensive Advancement of Beautiful China Construction was issued by the Central Committee of the Communist Party of China and the State Council clearly clarified the exploration and development of ESG assessment frameworks, which shows a strong will of establishing ESG Evaluation System with Chinese Characteristics.
-Despite the implementation of various ESG policies in recent year, there are still significant disagreement in ESG rating system.
-From the moment the ESG concept was introduced in China, multiple rating agents have independently published their ESG rating reports.
-However, these reports often display significant discrepancies in ratings, which undoubtedly has a substantial impact on the companies being evaluated.
-Examining the economic consequences of these ESG rating disagreements is of considerable value for both enterprises and regulatory bodies.
-The main recent studies are: Wang et al.
-(2024); Zhao et al.
-(2024) argue that disagreements in ESG ratings are significantly negatively correlated with stock returns; ESG rating disagreement increase analyst forecast errors(Liu et al., 2024); ESG rating disagreement drives firms towards green innovation (Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); Moreover, increased disagreement in ESG ratings leads to higher levels of real earnings management both in the current period and over the next one to two years (Li and Cheng, 2024); The firms with agreed high ESG rating can benefit through cost reduction and larger amount debt-financing(Guo et al., 2024); Besides,the disagreement in ESG ratings may poses a potential risk of stock price collapse(Luo et al., 2023); What’s more, ESG rating disagreement significantly increases audit fees due to information asymmetry(Ling et al., 2024).
-However, the focus of most recent studies has been less on exploring the impact of ESG rating disagreements on the overall development of a company.
-Therefore, we select TFP as the subject of our analysis in this study, as it serves as an indicator of a company’s high-quality development.
+Since 1987, the United Nations has championed sustainable development and introduced ESG (Environmental, Social, and Governance) in 2004.
+Over time, these principles have become key components of global investment strategies and corporate assessments.
+In recent years, China has embraced sustainable development through initiatives like the philosophy “green mountains and clear waters are as valuable as mountains of gold and silver.” On December 27, 2023, the Central Committee of the Communist Party of China and the State Council issued the Guidelines for the Comprehensive Advancement of Beautiful China Construction, which emphasized creating an ESG Evaluation System with Chinese Characteristics.
+Despite recent ESG policy efforts, significant disagreements persist within ESG rating systems.
+Since ESG ratings were introduced in China, multiple agencies have issued their own reports, often resulting in substantial discrepancies.
+Such differences have considerable implications for evaluated companies.
+Recent studies highlight the economic impacts of ESG rating disagreement: stock returns suffer (Wang et al., 2024; Zhao et al., 2024); analyst forecast errors rise (Liu et al., 2024); green innovation encouragement (Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); real earnings management increases in both the short and long term (Li and Cheng, 2024); firms with consistently high ratings benefit from lower costs and increased debt financing (Guo et al., 2024); disagreement poses a risk of stock price collapse (Luo et al., 2023); audit fees increase due to information asymmetry (Ling et al., 2024); and market pressures on managers intensify (Chen and Wang, 2025).
+However, most recent studies have not sufficiently examined ESG rating disagreement’s impact on a company’s overall development.
+To address this gap, we focus on TFP as an indicator of high-quality development in this study.
 2 Theoretical Analysis and Research Hypotheses
-According to information asymmetry theory, ESG rating disagreements can confuse stakeholders and contribute to the degree of information asymmetry between companies and investors.
-ESG ratings, as non-financial information, provide stakeholders with a more comprehensive insight into companies, thereby alleviating information asymmetry(Xiaohong and Zhenghan, 2023).
-Conversely, significant ESG rating disagreements can intensify information asymmetry, creating a noise effect that leads to investor misjudgments(Berg et al., 2022).
-Thereby, increasing financing constraints and ultimately resulting in a decline in TFP.
-Simultaneously, based on stakeholder theory, companies may take certain remedial measures in response to ESG rating disagreements to balance the interests of various stakeholders.
-Companies might intensify green innovation under the stimulus of ESG rating disagreements(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); However, as this initiative cannot cater to all stakeholders under noise effect caused by ESG rating disagreement, it may lead some groups to perceive "greenwashing" behaviors, which ultimately undermines the company’s financing efforts and results in a decline in TFP(Peng and Xie, 2024).
-Meanwhile, the divergence in ESG ratings may also drive companies to focus on improving their ESG recognition, potentially causing an imbalance between the interests of external investors and internal employees.
-This imbalance could lead to a human capital outflow, ultimately affecting the development of TFP.
-In summary, we propose the research hypothesis H1: ESG rating disagreement reduces corporate TFP.
+ESG rating disagreement can confuse stakeholders and increase information asymmetry between companies and investors.
+While ESG ratings help reduce information asymmetry by providing non-financial insights(Xiaohong and Zhenghan, 2023), disagreement can amplify asymmetry, creating noise that misleads investors(Berg et al., 2022), raises financing constraints, and ultimately lowers TFP.
+Stakeholder theory suggests that companies may respond to disagreement by pursuing green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a), but this can appear as “greenwashing” to some stakeholders, undermining financing and reducing TFP(Peng and Xie, 2024).
+Disagreement may also prompt firms to prioritize ESG recognition, causing tension between external investors and internal employees, potentially leading to a human capital outflow and further TFP decline.
+Hence, we propose the research hypothesis H1: ESG rating disagreement reduces corporate TFP.
 3 Data and Empirical Design
 3.1 Sample Selection
-We utilize financial data of A-share listed companies in China from 2015 to 2022111Prior 2015, the widely recognized local ESG rating agency in China was Huazheng., alongside ESG ratings from four agencies: Huazheng, Wind, SynTao Green Finance, and Susallwave.
-The data processing procedures follow these steps: (1) excluding companies in the financial and real estate industries; (2) excluding ST, *ST, and delisted companies; (3) excluding companies with only one ESG rating in a given year; (4) excluding companies with missing key indicators; and (5) applying Winsorization at the 1% level on both tails.
-All data is sourced from the CSMAR database, the CNRDS database, and the four rating agencies, resulting in 3656 firms and 13417 samples in baseline regression.
+We use financial data of A-share listed companies in China from 2015 to 2022111Prior 2015, the widely recognized local ESG rating agency in China was Huazheng., along with ESG ratings from Huazheng, Wind, SynTao Green Finance, and Susallwave.
+The data processing steps are as follows: (1) exclude companies in the financial and real estate sectors; (2) exclude ST, *ST, and delisted firms; (3) exclude companies with only one ESG rating in a year; (4) exclude firms with missing key indicators; (5) apply Winsorization at the 1% level on both tails.
+The data comes from the CSMAR and CNRDS databases and the four rating agencies, yielding 3656 firms and 13,417 samples in the baseline regression.
 3.2 Model Specification
@@
 The main variables are defined in Table 1.
-Table 1: Definition of Main Variables
-Symbol Source Description
-$TFP$ CSAMR Estimated using the LP, OP and GMM method
-$Dis$ ESG Rating Agencies Calculated based on the method described above
-$TobinQ$ CSAMR Tobin’s Q ratio
-$Top10$ CSAMR Shareholding of top 10 shareholders / total shares
-$ListAge$ CSAMR Natural logarithm of years since the firm’s IPO
-$Lev$ CSAMR Total liabilities at year-end / total assets at year-end
-$Size$ CSAMR Natural logarithm of total assets at year-end
-$Avg$ CSAMR Natural logarithm of operating income / number of employees
-$Der$ CSAMR Total liabilities at year-end / total equity at year-end
-$EI$ CNRDS Natural logarithm of the number of green patents obtained in a given year
-$FC$ CSAMR WW and KZ indices, the larger they are the more constrained the financing is
-$HC$ CSAMR Natural logarithm of number of graduate students and above and natural logarithm of number of undergraduate students and above
+[Insert Table 1 Here]
 4 Empirical Analysis
@@
 The results indicate that ESG rating disagreement reduces corporate TFP, providing preliminary support for the validity of Hypothesis H1.
-Table 2: Baseline Regression Results
-Variables TFP
-$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
-$Dis$0.0675*** -0.0536** -0.0536** -0.0447*** -0.0357** -0.0373**
-(-2.7452) (-2.1838) (-2.2098) (-2.9449) (-2.2503) (-2.2395)
-$TobinQ$0.0032 -0.0017 -0.0007
-(-0.5471) (-0.3953) (-0.1795)
-$Top10$0.0010 -0.0009 -0.0005
-(-0.9976) (-0.9496) (-0.4838)
-$ListAge$ 0.0989** 0.0855** 0.0160
-(2.4952) (2.3667) (0.4241)
-$Lev$0.1697 -0.1557 -0.1481
-(-1.5121) (-1.5361) (-1.4348)
-$Size$ 0.3737*** 0.2581*** 0.2245***
-(13.129) (14.214) (10.939)
-$Avg$ 0.2292*** 0.2695*** 0.2671***
-(2.7852) (2.8274) (2.7426)
-$Der$$0*$0190* -0.0211**
-(-2.0405) (-1.8342) (-1.9816)
-Controls NO NO NO YES YES YES
-Year YES YES YES YES YES YES
-Id YES YES YES YES YES YES
-Doubel Clustering YES YES YES YES YES YES
-$N$ 13417 13417 13417 13417 13417 13417
-$R^{2}$ 0.0015 0.0011 0.0011 0.1785 0.1670 0.1409
-  • Note: t-statistics are in parentheses. * p<0.1, ** p<0.05, *** p<0.01. The same applies below.
+[Insert Table 2 Here]
 4.2 Mechanism Analysis
@@
 To examine the role of green innovation in the relationship between ESG rating disagreement and TFP, an interaction term between green innovation and ESG rating disagreement is constructed, with other settings remaining consistent with Equation 1.
-Table 3 and figure1 reports the results.
+Table 3 and Figure 1 report the results.
 The results show that when there are no ESG rating discrepancies, green innovation promotes the improvement of TFP.
@@
 According to stakeholder theory, firm’s remedial green innovations may be seen as "greenwashing" when there is disagreement over ESG ratings, which will lead to higher costs of debt financing for corporations, ultimately bringing about a decline in TFP(Peng and Xie, 2024).
-Table 3: Green Innovation Mechanism
-Variables TFP
-$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
-$Dis$0.0264 -0.0205 -0.0199 -0.0167 -0.0130 -0.0136
-(-1.2045) (-1.0114) (-0.9590) (-0.8367) (-0.7014) (-0.7069)
-$EI$ 0.0412*** 0.0283*** 0.0253*** 0.0223*** 0.0142*** 0.0126***
-(8.4064) (6.2383) (5.4522) (4.9596) (3.3848) (2.8967)
-$Dis*EI$0.0366*** -0.0292*** -0.0296*** -0.0249** -0.0199** -0.0207**
-(-3.1706) (-2.7348) (-2.7184) (-2.3693) (-2.0401) (-2.0432)
-$TobinQ$0.0030 -0.0016 -0.0007
-(-0.7895) (-0.4611) (-0.1778)
-$Top10$$0$$0$0005
-(-1.7413) (-1.7832) (-0.8682)
-$ListAge$ 0.1012*** 0.0869*** 0.0172
-(3.4856) (3.2139) (0.6122)
-$Lev$0.1692*** -0.1550*** -0.1472***
-(-3.2834) (-3.2318) (-2.9594)
-$Size$ 0.3664*** 0.2537*** 0.2208***
-(28.243) (21.017) (17.632)
-$Avg$ 0.2297*** 0.2698*** 0.2673***
-(22.666) (28.608) (27.331)
-$Der$$0**$0190*** -0.0211***
-(-3.1573) (-3.0646) (-3.2805)
-Controls NO NO NO YES YES YES
-Year YES YES YES YES YES YES
-Id YES YES YES YES YES YES
-$N$ 13417 13417 13417 13417 13417 13417
-$R^{2}$ 0.0092 0.0052 0.0042 0.1806 0.1680 0.1417
-Figure 1: Marginal Effect
-To further investigate whether the moderating effect varies across levels of ESG rating disagreement, we conducted regressions separately for the upper and lower 50% quartiles of ESG rating disagreement, and the tabel 4 show that the moderating effect is significant only when the ESG rating disagreement is high, further supporting the notion that substantial ESG rating disagreement may lead to stakeholder distrust in firms’ green innovation.
-Table 4: Heterogeneity of Green Innovation Mechanism
-Variables TFP
-High $Dis$ Low $Dis$
-$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
-$Dis$ 0.0527 0.0616 0.0501 -0.2268*** -0.1532** -0.1554**
-(1.2360) (1.5631) (1.2368) (-2.8516) (-2.0471) (-1.9903)
-$EI$ 0.0364*** 0.0280*** 0.0230** 0.0092 0.0043 0.0041
-(3.1495) (2.6193) (2.0914) (1.4601) (0.7258) (0.6628)
-$Dis*EI$0.0522** -0.0461** -0.0381** 0.0433 0.0280 0.0260
-(-2.1556) (-2.0558) (-1.6525) (1.2832) (0.8798) (0.7855)
-$TobinQ$0.0009 -0.0012 0.0020 -0.0025 -0.0020 -0.0013
-(-0.1395) (0.1681) (0.3447) (-0.4475) (-0.3853) (-0.2318)
-$Top10$0.0008 -0.0012 -0.0005 -0.0001 0.0004 0.0008
-(-0.8389) (-1.3029) (-0.5405) (-0.1141) (0.4905) (0.8934)
-$ListAge$ 0.1137** 0.0833* 0.0257 0.0609 0.0529 -0.0332
-(2.2567) (1.7886) (0.5359) (1.4289) (1.3186) (-0.7922)
-$Lev$0.1140 -0.0997 -0.0976 -0.2786*** -0.2703*** -0.2712***
-(0.1716) (-1.2930) (-1.2306) (-3.2945) (-3.3962) (-3.2670)
-$Size$ 0.3607*** 0.2525*** 0.2173*** 0.3523*** 0.2355*** 0.2123***
-(16.816) (12.732) (10.655) (17.502) (12.434) (10.746)
-$Avg$ 0.2126** 0.2424*** 0.2447*** 0.2301*** 0.2706*** 0.2651***
-(13.198) (16.277) (15.983) (13.969) (17.451) (16.389)
-$Der$ 0.0224** -$0*$0226** 0.0046 0.0052 0.0036
-(-2.2377) (-2.2735) (-2.3751) (0.3638) (0.4404) (0.2910)
-Controls YES YES YES YES YES YES
-Year YES YES YES YES YES YES
-Id YES YES YES YES YES YES
-$N$ 6774 6774 6774 6643 6643 6643
-$R^{2}$ 0.1578 0.1434 0.1229 0.1717 0.1533 0.1310
-4.2.2 The Role of Financing Constraints
+[Insert Table 3 Here]
+[Insert Figure 1 Here]
+To further investigate whether the moderating effect varies across levels of ESG rating disagreement, we conducted regressions separately for the upper and lower 50% quartiles of ESG rating disagreement, and the Table 4 show that the moderating effect is significant only when the ESG rating disagreement is high, further supporting the notion that substantial ESG rating disagreement may lead to stakeholder distrust in firms’ green innovation.
+[Insert Table 4 Here]
+4.2.2 The Role of Financing Constraint
 Undoubtedly, ESG investment plays a crucial role in promoting ESG development, making it essential to examine the impact of ESG rating disagreement on corporate financing.
@@
 We only use independent variable to regress with finacing constraints, since numerous researches have demonstrated that financing constraints are detrimental to TFP improvement(Hopenhayn, 2014; Zhang et al., 2021; Piao et al., 2023).
-Table 5 reports the regression results. The $Disb$ is the variable by one period. So we concludes that increased ESG rating disagreement reduces corporate financing constraints, thereby lowering TFP.
-Table 5: Financing Constraints Mechanism
-Variables Financing Constraints
-$WW$ $WW$ $KZ$ $KZ$ $WW$ $WW$ $KZ$ $KZ$
-$Dis$ 0.0122*** 0.2166*** 0.0089*** 0.1815***
-(4.7408) (3.0974) (3.6730) (2.6598)
-$Disb$ 0.0122*** 0.2142*** 0.0074*** 0.1392**
-(4.7691) (3.0604) (3.3430) (2.2547)
-$TobinQ$0.0022*** -0.0006 0.2724*** 0.0237
-(-3.9269) (-0.9530) (19.528) (1.4433)
-$Top10$0.0007*** -0.0006*** 0.0011 -0.0061**
-(-8.8614) (-6.9840) (0.4783) (-2.3496)
-$ListAge$ 0.0204*** 0.0143*** 0.5530*** 1.2048***
-(6.1906) (2.9544) (6.5333) (9.5882)
-$Lev$ 0.0370*** 0.0359*** 6.7446*** 3.1453***
-(5.0820) (4.5416) (32.842) (14.109)
-$Size$0.0614*** -0.0455*** -0.4512*** -0.2274***
-(-36.832) (-22.973) (-9.5683) (-4.0781)
-$Avg$$0**$0113*** -0.5288*** -0.0107
-(-18.835) (-7.3282) (-13.531) (-0.2432)
-$Der$ 0.0030*** 0.0011 0.0191 -0.0027
-(3.4646) (1.2994) (0.6820) (-0.0931)
-Controls NO NO NO NO YES YES YES YES
-Year YES YES YES YES YES YES YES YES
-Id YES YES YES YES YES YES YES YES
-$N$ 10059 10059 13417 13417 10059 10059 13417 13417
-$R^{2}$ 0.0034 0.0034 0.0010 0.0010 0.2647 0.1221 0.2250 0.0517
-  • Note: Similar to the robustness tests, $Dis$ corresponds to regressions on ESG rating disagreement in year t and financing constraints in year t+1, while $Disb$ corresponds to regressions on ESG rating disagreement in year t-1 and financing constraints in year t.
+Table 5 reports the regression results.
+The $Disb$ is the variable by one period.
+So we concludes that increased ESG rating disagreement strengthens corporate financing constraints, thereby lowering TFP.
+[Insert Table 5 Here]
 4.2.3 The Role of Human Capital
@@
 Based on this evidence, we argue that Dis reduces TFP by impairing human capital.
-Table 6: Human Capital Mechanism
-Variables Human Capital
-$Master$ $Master$ $Undergrad$ $Undergrad$ $Master$ $Master$ $Undergrad$ $Undergrad$
-$Dis$0.0595*** -0.0526*** -0.0345** -0.0316**
-(-2.8695) (-3.1495) (-1.7625) (-2.0461)
-$Disb$0.0598*** -0.0528** -0.0322* -0.0248*
-(-2.8867) (-3.1608) (-1.8268) (-1.9302)
-$TobinQ$ 0.0197*** -0.0067 0.2724*** -0.0036
-(5.0228) (-1.4155) (19.528) (-0.9603)
-$Top10$ 0.0029*** 0.0008 0.0011 0.0003
-(4.4620) (1.0079) (0.4783) (0.4613)
-$ListAge$0.0010 0.1365*** 0.5530*** 0.1578***
-(-0.0413) (3.7156) (6.5333) (5.5267)
-$Lev$ 0.0536 -0.0620 6.7446*** 0.0894*
-(0.8913) (-0.9460) (32.842) (1.7679)
-$Size$ 0.6722*** 0.4376*** -0.4512*** 0.4410***
-(49.844) (27.508) (-9.5683) (35.103)
-$Avg$0.1597*** -0.0907*** -0.5288*** -0.0739***
-(-14.054) (-7.0579) (-13.531) (-7.5015)
-$Der$0.0256*** -0.0128 0.0191 -0.0184***
-(-3.1008) (-1.5259) (0.6820) (-2.8025)
-Controls NO NO NO NO YES YES YES YES
-Year YES YES YES YES YES YES YES YES
-Id YES YES YES YES YES YES YES YES
-$N$ 10461 10461 12595 12595 10461 10461 12595 12595
-$R^{2}$ 0.0011 0.0011 0.0010 0.0011 0.2790 0.1145 0.4112 0.1511
-  • Note: Similar to the robustness tests, $Dis$ corresponds to regressions on ESG rating disagreement in year t and human capital in year t+1, while $Disb$ corresponds to regressions on ESG rating disagreement in year t-1 and human capital in year t.
+[Insert Table 6 Here]
 4.3 Heterogeneity Analysis
 To further explore whether the dampening effect of ESG rating disagreement on TFP exhibits heterogeneity, this study conducts a subgroup regression analysis.
-Given that the sample size decreases after subgrouping, making the results more susceptible to the influence of outliers, the data was Winsorized at the 2% level on both ends before performing the subgroup regression and conducting a Chow Test on $Dis$.
+Given that the sample size decreases after subgrouping, making the results more susceptible to the influence of outliers, the data is Winsorized at the 2% level on both ends before performing the subgroup regression and conducting a Chow Test on $Dis$.
 Table 7 reports the results of the heterogeneity analysis.
@@
 Consequently, the negative impact of disagreements on TFP is more significant for high-tech firms.
-Table 7: Heterogeneity Analysis
-Variables TFP
-State-Owned Non-State-Owned Capital-Intensive Non-Capital-Intensive High-Pollution Low-Pollution High-Tech Non-High-Tech
-$TFP_{LP}$
-$Dis$0.0622** -$0$0322 -0.0462*** -0.0239 -0.0481*** -0.0425** -0.0043
-(-2.0516) (-1.8294) (-1.1009) (-2.8884) (-1.0401) (-2.7959) (-2.2163) (-0.1653)
-Chow Test 500.19*** 289.12*** 544.57*** 756.47***
-$R^{2}$ 0.1573 0.1886 0.1741 0.1817 0.2010 0.1754 0.1751 0.1570
-$TFP_{OP}$
-$Dis$0.0553* -0.0268 -0.0097 -0.0458*** -0.0070 -0.0493** -0.0365** -0.0048
-(-1.8608) (-1.5208) (-0.3229) (-2.6338) (-0.3149) (-2.4735) (-2.2763) (-0.1639)
-Chow Test 529.58*** 269.31*** 534.66*** 760.50***
-$R^{2}$ 0.1759 0.1593 0.1672 0.1663 0.1930 0.1595 0.1561 0.1526
-$TFP_{GMM}$
-$Dis$0.0618** -$0$0278 -0.0471*** -0.0134 -0.0519** -0.0413** -0.0122
-(-2.1381) (-1.6745) (-0.9906) (-2.6889) (-0.6185) (-2.5484) (-2.4165) (-0.4205)
-Chow Test 498.97*** 248.15*** 475.42*** 698.54***
-$R^{2}$ 0.1607 0.1254 0.1504 0.1377 0.1691 0.1321 0.1256 0.1315
-Controls YES YES YES YES YES YES YES YES
-Year YES YES YES YES YES YES YES YES
-Id YES YES YES YES YES YES YES YES
-Double Clustering YES YES YES YES YES YES YES YES
-$N$ 4416 9001 2342 11075 3900 9517 8454 4963
+[Insert Table 7 Here]
 5 Further Discussion
-In the above analysis we can get that ESG rating disagreement brings about a decrease in TFP, which is obtained based on linear regression.
+In the above analysis we can see that ESG rating disagreement brings about a decrease in TFP, which is obtained based on linear regression.
 In order to further explore the relationship between ESG rating divergence and TFP under the nonlinear assumption, we consider a regression analysis using a machine learning approach that can be interpreted with SHapley Additive exPlanations(SHAP) introduced by Scott et al.
@@
 (2024), we use the XGBoost machine learning model introduced by Chen and Guestrin (2016) to predict $TFP_{LP}$.
-XGBoost is an ensemble learner based on gradient boosted trees (GBDT) that has demonstrated excellent performance in many applications, even outperforming deep neural networks in some tasks.
+XGBoost is an ensemble learner based on gradient boosted trees (GBDT) that has demonstrated excellent performance in many applications, even outperforming deep neural networks in tabular data(Grinsztajn et al., 2022).
 In this paper, the Optuna library developed by Akiba et al.
@@
 This improvement underscores the predictive value of ESG ratings in modeling TFP.
-Table 8: Performance Metrics for Training and Test Sets
-Metric Training Set 1 Training Set 2 Training Set 3 Training Set 4 Test Set 1 Test Set 2 Test Set 3 Test Set 4
-MSE 0.2609 0.2274 0.2818 0.4714 0.3492 0.3300 0.3835 0.4728
-RMSE 0.5108 0.4769 0.5309 0.6866 0.5909 0.5744 0.6192 0.6876
-MAE 0.3995 0.3729 0.4145 0.5383 0.4589 0.4523 0.4818 0.5415
-MAPE 4.7786% 4.4569% 4.9457% 6.4047% 5.4649% 5.3644% 5.7293% 6.4399%
-R2 0.7685 0.7969 0.7500 0.5794 0.7045 0.7176 0.6755 0.5955
-  • Note: Set1 is the XGBoost performance of $Dis$ in the same period as the TFP, Set2 is the XGBoost performance of $Dis$ in the preceding period, Set3 is the XGBoost performance excluding $Dis$, and Set4 is the model performance of the linear model (no fixed effects) and $Dis$ in the preceding period.
-To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in our predictive task, we applied SHAP to interpret the results of the XGBoost model.
+[Insert Table 8 Here]
+To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in our predictive task, we use SHAP to interpret the results of the XGBoost model.
 SHAP values are grounded in cooperative game theory and provide a way to understand each feature’s contribution to the model output by treating features as "players" in a game.
@@
 $\phi_{i}=\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!\cdot(|N|-|S|-1)!}{|N|!}\left[f(S\cup\{i\})-f(S)\right]$ (3)
-In this equation:
-  • $N$ represents the set of all features, while $S$ is a subset of $N$ that does not include feature $i$.
-  • $|S|$ denotes the size of subset $S$.
-  • $f(S)$ is the model output based on the subset of features $S$.
-  • $f(S\cup\{i\})$ is the output when feature $i$ is added to $S$.
+where $N$ represents the set of all features, $S\subseteq N\setminus\{i\}$ is a subset excluding feature $i$, $|S|$ denotes the cardinality of subset $S$, $f(S)$ indicates the model output using feature subset $S$, and $f(S\cup\{i\})$ represents the output when feature $i$ is added to $S$.
 The term $\frac{|S|!\cdot(|N|-|S|-1)!}{|N|!}$ in the equation represents a weighting factor that accounts for all possible orders in which features can be added to the model.
@@
 In this study, based on the previous training results, we chose the model with one-period lagged ESG rating disagreement for our SHAP analysis, and the SHAP values were calculated for the test set.
-Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting TFP [^3] The importance of each feature is calculated as the percentage of the average of the absolute SHAP values from each to the mean of the absolute SHAP values of all features; in short, a higher percentage means that the feature is more important among all features..
-Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s TFP, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task [^4] The beeswarm plot shows the effect of different values of feature on TFP, the color of the dots in the plot shifts toward red for larger eigenvalues and toward blue for smaller eigenvalues and points positioned on the horizontal axis with values less than 0 indicate a negative effect on the dependent variable in the model, while points with values greater than 0 signify a positive effect..
+Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting TFP.
+Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s TFP, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task.
 The reason for this phenomenon is that the investment market is more tolerant of small ESG rating divergence, and listed companies will not face significant financing pressure due to small divergence.
-Figure 2: Feature Importance
-Figure 3: SHAP Beeswarm Plot
+[Insert Figure 2 Here]
+[Insert Figure 3 Here]
 6 Conclusion and Implication
-ESG rating disagreements present significant challenges to the improvement of corporate TFP, underscoring the urgent need for firms to address this issue.
-Mechanism analysis reveals that green innovation can exacerbate the negative effects of ESG rating disagreements on TFP.
-Additionally, these disagreements hinder TFP growth by intensifying financing constraints and weakening human capital.
+ESG rating disagreement presents significant challenges to the improvement of corporate TFP, underscoring the urgent need for firms to address this issue.
+Mechanism analysis reveals that green innovation can exacerbate the negative effects of ESG rating disagreement on TFP.
+Additionally, these disagreement hinder TFP growth by intensifying financing constraints and weakening human capital.
 To mitigate these impacts, companies should focus on improving the quality of ESG information disclosure, carefully navigate the trade-offs between green innovation, financing needs, and strive to balance the interests of diverse stakeholders.
 Regulatory authorities, in turn, should strengthen the scrutiny of green innovation patents to ensure their quality and prevent superficial compliance efforts.
-Heterogeneity analysis demonstrates that the negative effects of ESG rating disagreements on TFP are more pronounced in state-owned enterprises, non-capital-intensive firms, low-pollution industries, and high-tech enterprises.
+Heterogeneity analysis demonstrates that the negative effects of ESG rating disagreement on TFP are more pronounced in state-owned enterprises, non-capital-intensive firms, low-pollution industries, and high-tech enterprises.
 These entities should pay closer attention to ESG rating discrepancies, while regulators should provide targeted support to low-pollution and high-tech firms to foster sustainable development and technological innovation.
-Moreover, ESG rating disagreements serve as a critical predictor of TFP, with SHAP values highlighting their substantial and ongoing negative impact.
+Moreover, ESG rating disagreement serve as a critical predictor of TFP, with SHAP values highlighting their substantial and ongoing negative impact.
 This indicator can be leveraged by listed companies and regulatory agencies as a monitoring and early warning tool to identify potential risks to TFP and implement proactive measures.
-Appendix A Appendix
+We believe it would be worthwhile to further investigate how green innovation amplifies the negative impact of ESG rating divergence on TFP, as well as to integrate SHAP with microeconometric methods for more robust empirical analysis.
+Appendix A Online Appendix
 A.1 Descriptive Statistics
@@
 $IV$ 0.082 0.067 0.059 -0.974 -0.008 0.087 -0.024 -0.038 0.082 0.027 -0.046 0.014 1.000
-A.3 Robustness Tests and Endogeneity Treatment
+  • Note: This is the matrix of linear correlation coefficients for all variables.
+A.3 Robustness Tests and Endogeneity Treatments
 To ensure the robustness of the baseline regression results, the following robustness tests were conducted.
@@
 $R^{2}$ 0.0015 0.0011 0.0011 0.7027 0.8496 0.7734
+  • Note: This is the result of only front-loading $Dis$ compared to $TFP$ by one period.
+In this regression result, $Dis$ denotes ESG Rating Disagreement of $t-1$ year and $TFP$ denotes total factor productivity of $t$ year.
+Unless otherwise noted, $Dis$ and $TFP$ are consistent with the baseline regression settings.
 Table 12: Robustness Tests (II)
@@
 $R^{2}$ 0.0014 0.0011 0.0010 0.1793 0.1670 0.1406
+  • Note: This is the result of the regression with the addition of control variables, and the other settings are consistent with the baseline regression.
+$N.TFP_{province}$ represents the level of total factor productivity at the provincial level.
 Table 13: Robustness Tests (III)
@@
 R2 0.9517 0.1794 0.1671 0.1408
-  • Note: TFP is t+1 period, just like the baseline regression setup.
+  • Note: This is the result of an instrumental variable test where $TFP$ is the t+1 period, just like the baseline regression.
 Table 14: Robustness Tests (IV)
@@
 R2 0.0028 0.0027 0.0024 0.1739 0.1628 0.1455
+  • Note: This is the result after PSM, matched on a 1:1 basis, with other settings not differing from the baseline regression.
 A.4 Optuna Parameter Range
@@
 seed 42
+  • Note: This table shows the range of hyperparameters Optuna searches for, automatically selecting the optimal hyperparameters for the final model configuration.